[Search for users] [Overall Top Noters] [List of all Conferences] [Download this site]

Conference asdg::cmp

Title:CMP
Notice:Welcome to the CMP notesfile
Moderator:FABSIX::B_TOWERLLE
Created:Tue Oct 31 1995
Last Modified:Fri Jun 06 1997
Last Successful Update:Fri Jun 06 1997
Number of topics:94
Total number of notes:6381

21.0. "CMP.B3 EE/PE PASSDOWNS" by STRATA::TDYER () Tue Mar 26 1996 08:08

    
    THIS NOTE IS FOR CMP.B3 EQUIPMENT PASSDOWNS
T.RTitleUserPersonal
Name
DateLines
21.1INSTALL CONTINUESSTRATA::WESTPHALSat May 04 1996 20:157
    
CMP.B3 - Reps continue with there installation checks and alignments.  The down
	 force for the carriers and the conditioner were calibrated.  Final
	 platen run out was noted.  Pressures and flows were also checked
	 and corrected as needed.  A pad was also put on as requested by
	 SPEEDFAM.

21.2Sunday (5/5) Installation UpdateSUBPAC::LANDRYSun May 05 1996 20:4332
    Today, the reps showed up to continue on with the start-up.
    Things accomplished today were:
    
    *  ALL DI water sprays were set up.  While setting up the sprays, the
       system periodically would alarm for LOW DI WATER FLOW.  There is a
       big pressure drop (~8 PSI) occassionally.  Turning up the pressure
       only created a different problem, which are that some of the DI
       solenoid valves remain ON when they should be shutting off.  Lower-
       ing the pressure back down allows them to shut off properly.  There
       is a possibility that the incoming water loop may need to supply a
       larger volume (not pressure) to the system.  It will be further
       looked at in the morning.
    
    *  The scrubber was occassionally timing out.  The physical movement
       was not acceptable.  The rails were fine.  The speed controlling
       valve was adjusted, but good, consistant results were never
       achieved.  The incoming air to the respective speed valve was in-
       creased to 60 PSI.  We may change the valve in the near future.
    
    *  As of 7:20 PM, the reps have run 100 wafers (NO_TOUCH program) very
       successfully.  More would have been run, but there was a waiting
       line forming at the OnTrak scrubbers.  They plan to run more before
       leaving for the day.
    
    *  Another "potential" check they may attempt (time permitting) is to
       calibrate the slurry flows.  They are currently running on the high
       side.
    
    *  No carry over work need be handed off to B shift tonight.  Reps are
       on or ahead of schedule.
    
    						A shift
21.3Slurry leak FABSIX::R_GEEMon May 06 1996 08:145
	The fitting coming out of the #2 slurry shutoff valve was
	leaking, repaired the leak and verified. The system is in
	wet mode.

					B shift
21.4di pressureSTRATA::RIDLONEliminate the obviousTue May 07 1996 02:3114
    A-shift mentioned having a problem with the DI low pressure alarm. From
    my understanding it happened most often when the elevators were
    filling. Fiddled around with the system and the only thing out of the
    ordinary that was noticed was that the elevator di bypasses were
    closed. These were opened up and then DI regulator readjusted up to
    24psi inside the machine. Also adjusted the di return regulator thats
    in the fluids cabinet so that the incoming DI regulator could be
    adjusted to a pressure of 30psi (gauge outside of machine)  which 
    mimics the pressures on b2 and b1. With these things adjusted the DI
    pressure holds fairly steady at 24-25psi and the lowest we saw it go
    when the elevators were filling was 16psi. 
    
    				
    						B-shift
21.5crash, due to pad issueSUBPAC::LANDRYTue May 07 1996 20:3630
      Today, final adjustments were made to the system.  More NO_TOUCH
      runs were made with good results.  Mid Afternoon, the reps decided
      to run some wafers touching down on the pad at about 80 lbs Down
      force.  A couple of table rotations into the run and CRASH.  Three
      carriers and the pad were trashed.  Upon removing the polishing
      pad, it was discovered that there was a piece (~ 3" X 10") of 
      plastic (the material that is on the sticky side of the new polishing
      pads) was still attached to the underside of the pad.  The gentle
      crackling sound of an air bubble (or the material on the underside
      of the pad) was heard right before the crash.  On another side note,
      the pad that was being used had dried out the previous day when the 
      power to the system off.  So, were chalking this crash up to the pad.

      The mild mess was cleaned up.  Three new carriers were installed.
      The system was left in WET MODE for the evening.

      The 500 wafer mechanical acceptance run is being planned for tomorrow
      morning, utilizing the prescribed programs layed out for the test.
      The two pads will be laid down in the morning (or B shift can do 
      this if they desire tonight).  Please keep in mind that there is 
      only ONE secondary table pad in-house.  It is in the draw second from 
      the bottom of the pad cabinet.  It is clearly labeled on the back side 
      to be used on B3 ONLY.  More pads are due in Friday (per my discussion
      with Linda Boisoneault).  If the primary pad is installed tonight, it
      can be conditioned as well, using the standard condition procedure.

      						Doc

 
21.6pads replacesFABSIX::R_GEEWed May 08 1996 07:374
	Placed a pad on both the primary and secondary table. Conditioner
	the primary pad, the system is in wet mode.

						B Shift
21.7Acceptance preparation day!SUBPAC::LANDRYWed May 08 1996 20:3745
      Well, we had a tentative plan to attempt the marathon run 
      today, but that was postponed pending further exercising of
      the system.  This due to the crash which occurred on 
      Tuesday afternoon.  Also, Roe and Hershel went to Fab 6
      protocol training in the morning. 

      What transpired today was actual testing of touchdown onto
      the Primary polish pad (and a couple of touchdowns on the 
      secondary pad).  All the recipes that are scheduled to run
      during the mechanical acceptance were tested briefly today
      (one set of five wafers on each recipe).  There were NO
      crashes, failures or assists.  The system made some noise
      (the noise that occurs when excessive vibration occurs) early
      in each of the runs, but the noise settled out a few 
      oscillations into each run.  

      There is no further work needed to be done by the oncoming shift
      tonight.  Please leave the system in WET MODE.

      Some noteables before the marathon run:

      *  The recipes on B3 have the slurry being dispensed from one
         pump at 1000 ML.  The existing tools flow at 551 ML, with both
         pumps on (or a total of 1102 ML).

      *  The recipes have been edited today to remove the second table
         polish.  This is due to the fact that this remains the one
         and only secondary polish pad in-house.  Please assure that
         the step is edited prior to the acceptance run.  ALL marathon
         recipes should be checked for this step.

      *  NO wafers reserved for the marathon have been used yet.  They
         all remain on the shelf next to the system.

      *  The pad was just put down Tuesday night/Wednesday morning.
         There are not many runs on it yet.  The secondary pad has
         only seen a couple of wafers.  (the ordered pads are due in 
         on Friday).

      *  The DI water situation seems resolved since B shift by-pass
         adjustments/regulator reset up.

      *  The acceptance run is scheduled to start Thursday morning.

      						A Shift
21.8start upSTRATA::TDYERFri May 10 1996 16:5041
    +---+---+---+---+---+---+---+
    | d | i | g | i | t | a | l |	  
    +---+---+---+---+---+---+tm-+
        Semiconductor
					I N T E R O F F I C E   M E M O

     To: CMP.dis		               Date: May 9th 1996
         		                       From: Timothy P. Dyer
                                               Dept: Engineering
     cc: Dennis Goodwin 	               Ext:  5072  
         		                       Node: Scoman::Tdyer
         
                            
                              
CMP.B3 - INSTALLATION and EQUIPMENT START UP are complete. We have started the
         EQUIPMENT ACCEPTANCE portion (500 wafer marathon). The marathon will
         continue into the night until the SF reps leave and will resume in 
         the morning. One DEC EEG tech must work with the reps to monitor 
         assist and or failures. I have allowed for 3 assist and no failures. 
         If a wafer STICKS to the pad during LIFT OFF (only) this can be 
         counted as an assist. Currently we have ONE assist logged, (wafer 
         stuck at lift off), this allows for TWO more possible assist. 
       
         wafer count	assist		failures    recipe     problem
       
            50		  0		   0        IO_MQC, 
           100		  0		   0        IO_bin1    
	   150            1                         IO_bin1    cassette 
           200            0                0        IO_bin1 
           250            0                0        IO_bin1 
           300                                      IO_bin1 
           350                                      IO_bin1 
           400                                      IO_bin1 
           450                                      IO_bin1 
           500	                                    IO_bin1,2,3,io_410 

  There is only ONE 2nd table pad for B3, it is on the tool. More pads are
  due in tomorrow.

   
                                            
21.9updateSTRATA::SWENSONSat May 11 1996 09:2749
    
              <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]CMP.NOTE;1 >>>
                                    -< CMP >-
================================================================================
Note 21.8                     CMP.B3 EEG PASSDOWNS                        8 of 8
STRATA::TDYER                                        41 lines  10-MAY-1996 15:50
                                 -< start up >-
--------------------------------------------------------------------------------
    +---+---+---+---+---+---+---+
    | d | i | g | i | t | a | l |	  
    +---+---+---+---+---+---+tm-+
        Semiconductor
					I N T E R O F F I C E   M E M O

     To: CMP.dis		               Date: May 9th 1996
         		                       From: Timothy P. Dyer
                                               Dept: Engineering
     cc: Dennis Goodwin 	               Ext:  5072  
         		                       Node: Scoman::Tdyer
         
                            
                              
CMP.B3 - INSTALLATION and EQUIPMENT START UP are complete. We have started the
         EQUIPMENT ACCEPTANCE portion (500 wafer marathon). The marathon will
         continue into the night until the SF reps leave and will resume in 
         the morning. One DEC EEG tech must work with the reps to monitor 
         assist and or failures. I have allowed for 3 assist and no failures. 
         If a wafer STICKS to the pad during LIFT OFF (only) this can be 
         counted as an assist. Currently we have ONE assist logged, (wafer 
         stuck at lift off), this allows for TWO more possible assist. 
       
         wafer count	assist		failures    recipe     problem
       
            50		  0		   0        IO_MQC, 
           100		  0		   0        IO_bin1    
	   150            1                         IO_bin1    cassette 
           200            0                0        IO_bin1 
           250            0                0        IO_bin1 
           300            0                0        IO_bin1 
           350                                      IO_bin1 
           400                                      IO_bin1 
           450                                      IO_bin1 
           500	                                    IO_bin1,2,3,io_410 

  There is only ONE 2nd table pad for B3, it is on the tool. More pads are
  due in tomorrow.

   
                                            
21.10additional update STRATA::SWENSONSat May 11 1996 09:3049
              <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]CMP.NOTE;1 >>>
                                    -< CMP >-
================================================================================
Note 21.8                     CMP.B3 EEG PASSDOWNS                        8 of 8
STRATA::TDYER                                        41 lines  10-MAY-1996 15:50
                                 -< start up >-
--------------------------------------------------------------------------------
    +---+---+---+---+---+---+---+
    | d | i | g | i | t | a | l |	  
    +---+---+---+---+---+---+tm-+
        Semiconductor
					I N T E R O F F I C E   M E M O

     To: CMP.dis		               Date: May 9th 1996
         		                       From: Timothy P. Dyer
                                               Dept: Engineering
     cc: Dennis Goodwin 	               Ext:  5072  
         		                       Node: Scoman::Tdyer
         
                            
                              
CMP.B3 - INSTALLATION and EQUIPMENT START UP are complete. We have started the
         EQUIPMENT ACCEPTANCE portion (500 wafer marathon). The marathon will
         continue into the night until the SF reps leave and will resume in 
         the morning. One DEC EEG tech must work with the reps to monitor 
         assist and or failures. I have allowed for 3 assist and no failures. 
         If a wafer STICKS to the pad during LIFT OFF (only) this can be 
         counted as an assist. Currently we have ONE assist logged, (wafer 
         stuck at lift off), this allows for TWO more possible assist. 
       
         wafer count	assist		failures    recipe     problem
       
            50		  0		   0        IO_MQC, 
           100		  0		   0        IO_bin1    
	   150            1                         IO_bin1    cassette 
           200            0                0        IO_bin1 
           250            0                0        IO_bin1 
           300            0                0        IO_bin1 
           350                                      IO_bin1 
           400                                      IO_bin1 
           450                                      IO_bin1 
           500	                                    IO_bin1,2,3,io_410 

  There is only ONE 2nd table pad for B3, it is on the tool. More pads are
  due in tomorrow.
  
    Speedfam rep will be in this a.m. 5/11 to complete marathon.
   
                                            
21.11Marathon Run Results/Adjustments/Ready for MQC'sASDG::POIRIERSat May 11 1996 20:5624
                                                         
    MARATHON RUN:
    ------------
    The 500 wafer marathon run was sucessfully completed today.
    2 assists were logged during the run. Both assists were 
    with the transfer paddle. 
    
    After the run was completed, the Speedfam rep found the a few
    problems with he transfer paddle which he corrected.
    
    a) The cylinder was in need of lubrication.
    b) The rail assembly was out of adjustment.
    c) Loose bolts were found in the assembly.
    
    The pre-aligner flow controllers were adjusted to ensure a 
    slow and steady clamping motion in the assembly. The fingers
    were slamming into the wafer prior to the adjustment.
    
    
    MQC's:
    -----
    The system is now ready for MQC's on B shift tonight.
    
    
21.12STRATA::SWENSONSat May 11 1996 21:0550
              <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]CMP.NOTE;1 >>>
                                    -< CMP >-
================================================================================
Note 21.8                     CMP.B3 EEG PASSDOWNS                        8 of 8
STRATA::TDYER                                        41 lines  10-MAY-1996 15:50
                                 -< start up >-
--------------------------------------------------------------------------------
    +---+---+---+---+---+---+---+
    | d | i | g | i | t | a | l |	  
    +---+---+---+---+---+---+tm-+
        Semiconductor
					I N T E R O F F I C E   M E M O

     To: CMP.dis		               Date: May 9th 1996
         		                       From: Timothy P. Dyer
                                               Dept: Engineering
     cc: Dennis Goodwin 	               Ext:  5072  
         		                       Node: Scoman::Tdyer
         
                            
                              
CMP.B3 - INSTALLATION and EQUIPMENT START UP are complete. We have started the
         EQUIPMENT ACCEPTANCE portion (500 wafer marathon). The marathon will
         continue into the night until the SF reps leave and will resume in 
         the morning. One DEC EEG tech must work with the reps to monitor 
         assist and or failures. I have allowed for 3 assist and no failures. 
         If a wafer STICKS to the pad during LIFT OFF (only) this can be 
         counted as an assist. Currently we have ONE assist logged, (wafer 
         stuck at lift off), this allows for TWO more possible assist. 
       
         wafer count	assist		failures    recipe     problem
       
            50		  0		   0        IO_MQC, 
           100		  0		   0        IO_bin1    
	   150            1                         IO_bin1    cassette 
           200            0                0        IO_bin1 
           250            0                0        IO_bin1 
           300            0                0        IO_bin1 
           350            0                1        IO_bin1    PADDLE UP 
           400            0                0        IO_bin1 
           450            0                0        IO_bin1 
           500	          1                0        IO_bin1,2,3,io_410 
                                                               paddle up
  There is only ONE 2nd table pad for B3, it is on the tool. More pads are
  due in tomorrow.

   
   After the 500 wafers were run the paddle assy was taken apart,cleaned and
   reassembled with 25 wafers cycled without reoccurance of the paddle time 
   out.                  
21.13mqc data for 5-13-96FABSIX::J_ROAMon May 13 1996 21:0026
    Mqc's performed today May 13/96
    
    Head  Mean      Unif %     R/R
      1   5890	    4.258      2945
      2	  5616	    3.655      2808
      3   5852	    2.953      2926
      4   5292	    4.378      3051
    
    Particles before
    
    Head  bin7	 total   scratch lenghth
     
      1   15-    39       0
      2   3      24       0
      3   2      19       0
      4   12     42       0
      5   3      21       0
    
    particles after
    
      1   188   575       0
      2   284   885     10.7
      3   133   411      1.37
      4   431  1452      1.36
      5   242   775      1.15
    
21.14Transfer paddle retaughtSTRATA::FINNMon May 13 1996 21:118

	Jim Sadin tried to run wafers to check Polish rates. Paddle did not 
rise up before indexing to the left and wedged itself into Teflon stanchion.
Reps retaught "home" and travel positions. Wafer were then completed.


					a shift
21.15passdownSTRATA::FINNTue May 14 1996 21:377
    
    	Process acceptance has begun. Changed pad and carriers per Frank
        before acceptance began. Since we are short on carriers, the ones
        that were removed are in water in the back room, Please keep them
        wet, they can be used in a pinch. They were in spec.
    
                                                      A shift
21.16Lockedout/Taggedout for the night!SUBPAC::LANDRYMon May 20 1996 20:2918
    The system is currently in a LOCKOUT/TAGOUT state.
    
    This is due to the removal of an Infranor.  "Said" Infranor has been
    installed on B2.  The Index Table on B2 stopped.  It was discovered
    that the problem was a blown Infranor (NO power within it, not power
    supply related).  Since we have none in stock, Speedfam officials
    authourized the removal from B3 to sustain B2 up status.  The reps
    have requested a new Infranor.  Should be in tomorrow morning.
    
    So, the system will remain locked out.  The CDA was shut off to allow
    the heads/carriers to lower into the load cups of the Index Table to
    retain a proper wet status.  The lockout is inplace, with Speedfam reps
    holding the key.  The water guns have been rigged to supply water to
    the Index Table and Polish Pad.  Please refrain from touching all
    setups until the new Infranor arrives.  Thanks in advance for your
    cooperation in this trying matter.
    
    					A Shift / Speedy Guys
21.17New trap intalled on vac pumpSTRATA::RIDLONEliminate the obviousTue May 21 1996 05:166
    New filter traps were found in 94 stock and installed onto B2 and B1.
    Replacement filter elements were also found and are in the top of the
    tool box with two more traps. 
    
    
    					B-shift
21.18B3's big adventureSUBPAC::CMPTue Jun 04 1996 17:384
    we have begun the installation of the motor on tool 024 (B3) we are
    half way completed with the motor pulled this morning and the new one
    replaced this afternoon. we hope to have it completed by thursday.
    
21.19doors derailedFABSIX::B_FINNMon Jun 10 1996 20:388

	3 of four sliding doors on b3 needed to be aligned to track, 
two were actually removed and replaced. Two bottom guide rail's were 
shimmed due to low spots.


					A shift 
21.20boot errorsSTRATA::FINNTue Jun 11 1996 21:2517
  	 Eng. tried to start a polish cycle. System had unknow wafers on 
carrier. He removed the mystery wafers and again tried to start cycle. This 
time the multihead traveled out to the table in a lost attempt top polish.
He EMO'ed the system and then started the system back up.  The system began to 
power up, but then shut down completely and would not boot up again. At 
this point I was called. Checked EMO and door panels then found breaker 
off for vacuum pump and CPU controller. Reset both and system powered up past 
K2, but software did not boot nor did the CRT come on. Rear CRT was also dead.
CPU has failure as well as all Digital and Analog I/O boards (dio/aio) the 
serial comm card (board 5)is working as well as all Dcx cards. Reset CPU with
no luck.. Traced power via schematic through system seems OK as does all 
power supplies. 5 and 24 volts is making it to the i/o's ... 
Looks like the CPU. Kevin has come back to the plant to work on it.

					A shift

21.21boot problem/no screenFABSIX::L_SHKOLNIKOVTue Jun 11 1996 22:4916
    arrived at digital at 20:00hr received pass down from Brendon. At tool b3
    observed what was happening and this is the course of action.
    1) powered tool down.
    2)reset every ckt brkr manually
    3)reset main power switch
    4) powered tool up to let it cycle, tool appeared to cycle but no
    screen on front
    5)reset cpu bd all bds cycled
    6) removed video splitter cable and reset cpu again, complete tool
    cycled up.
    7)powered off and hard cycled system from scratch and tool booted
    8) re-installed all carriers and tool is in wet mode
    will check video cable tomorrow
    KEVIN    


21.22motion problems w/polish tableSUBPAC::CMPWed Jun 12 1996 16:197
    We corrected a problem with jerky rotation on the polish table. We had
    looked at the belt tension being possibly to tight and we loosened the
    tension only to find the problem still there to a lesser degree, next
    we checked the wiring to the motor and discovered that 2 of the
    windings to the motor located in the junction box had become
    disconnected, we reconnected the wires and tested with no more jerky
    motion The tool is now fit for use.
21.23upgrade completionSUBPAC::CMPMon Jun 17 1996 14:134
    The upgrade for the bus bar and new style cables for the infranors are
    now completed, also done as part of the upgrade was voltage adjustments
    for the infranors and a software change on dcx.cmp (sets operating-
    parameters). Tool was tested and now ready for use.
21.24carrier at polish reed switchesFABSIX::R_GEETue Jun 18 1996 08:1021
	The system ran good for the first 3 hours of our shift.
	A very annoying error started appearing "Waiting for carriers 
	down" during polish and at the secondary pad, but infact they 
	all were down. It would display the error on the touch screen 
	with no buzzer and keep polishing until the STOP button was touched.
	When the STOP button was touched the carriers would raise and 
        the system would halt. 
	It wouldn't allow us to exit the process screen or press any 
	button but STOP, when you hit STOP twice the system would complete 
	the next function. For example, press STOP twice the MTHA moved
	to the polish table, press STOP two more times , unload flipper 
	would move to the unload cup, press STOP two more times, unload 
	flipper to water drain position. This would go on until all 5 
	wafer in the index table were put into the exit cassette. The 
	system would then allow you to exit the process screen, fun stuff.
        	
	We found that the reed switches for heads #3, #4, & #5 at polish
	table and at wafer clean positions were intermittently not being 
	made. Adjusted the reed switches and ran test wafers. 

					B Shift
21.25elevator 1 home offSTRATA::RIDLONEliminate the obviousWed Jun 19 1996 05:3410
    Cassette 1 was double loading in slot 1 only, the rest of the cassette
    would load fine. Found that the first wafer was hitting the cassette.
    Recalibrated the home for elevator 1 so that the first wafer would load
    smoothly. Had to decrease the counts by about 3000 before the wafer
    would load smoothly which led us to check the elevator 1's sensor. It
    was loose. After it was tightened we recalibrated elevator 1's home and
    then tested a cassette load of wafers with no problems encountered. 
    
    
					b-shift
21.26cabling correctedSUBPAC::CMPWed Jun 19 1996 11:126
    The cabling for the keyboard is now plugged back in so the front and
    rear of the tool can now be accessed when needed.
    
    
    
                                                  Rich (Mr. Goodwrench)
21.27Slurry flow problems and touch screen issueSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringFri Jun 21 1996 08:3322
    
            Came to have both slurry pumps not working.  Went through 
    schematics and found loose connectors on sb#12 and tb#4.  Both 
    slurry pumps now working.  The slurry flow was now checked.  
    slurry pump #1 was flowing fine.  Pump #2 was checked and only
    DI water came out.  We found the Slurry Pump #2 DI rinse flow
    control all the open causing only DI water to flow through
    the tool.  The flow control was closed several turns to match
    #1.  The flow was tested and the flow amount was ok and it was
    only slurry.  The rinse was the turned on and it purged the 
    lines clear.


       Something else was happening this morning.  After the tool was 
    powered down due to frozen screen we powered it up and got the 
    error message ellodev not installed divide overflow  - so we removed 
    all communication and video cable which goes to rear of the tool 
    and disconnect second monitor.  The tool powered up fine.

                                            D shift


21.28Slurry Delivery Trouble / Timeout Errors - Carriers at ScrubASDG::POIRIERFri Jun 21 1996 20:3029
    
    * Slurry deliverly / sensor errors:
      --------------------------------
    
      System timed out for "No Slurry Flow". There was an air pocket in the
      top 8" of supply tubing just under the sensor in the Fluids Cabinet.
      Slurry was flowed manually and the air pocket disappeared only to
      return 15 min later. Ashland was called and told us that they had
      just changed a slurry filter. This activity evidently causes  air
      pockets to flow through the supply system and end up in the Fluids
      Cabinet at the Slurry Sensor location. THis causes errors!
    
      a) Ashland has agreed to notify the CMP area whenever filter changes 
         are done. 
    
      b) We are considering a new location for the slurry sensor to avoid this
         problem in the future.
    
    
      Timeout Error for "Carrier At Scrub":
      -------------------------------------
      The system errored for this problem today. The system displayed
      "Press any button to continue". Earl tried this and it timed out
      again. The tool was EMO'd, rebooted and 5 wafers were sucessfully
      run for verification.
    
    
     
      event.   
21.29high particles and uniformityFABSIX::R_GEEMon Jun 24 1996 08:3611
	The system failed mqc's for high uniformity and particles.

	Changed the two pads and ran breakins then mqc's. The mqc's
	failed for particles and head #4 had high uniformity. 
	We changed carrier #4 and also the scrub brushes. Ran mqc's
	again, the uniformity failed with all heads around 11%,
	particles are still out of spec also at 300. The ID/OD will
	need to be adjusted and hopefully running more wafers will
	bring the particles down. 

					B Shift
21.30Head #3 rotation motor controllerSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringThu Jun 27 1996 05:2013
      B3 was powered down for the weekly PM to check
   for leaks.  After everything was cleaned up, the tool
   was powered back up but head #3 wasn't rotating.  We
   lost the Infranor apparently from either the EMO or
   power-up.  The tool was off for a good 15-20 minutes.
   There are no rotation Infranors in stock.  Kev from
   SPEEDFAM has one and will be bringing it into the fab
   at the change of shift.  The tool has a WARNING on the
   tool for head #3 but the tool can be run if needed.

                                     D shift 

21.31faulty wafer loss detectionSTRATA::WESTPHALFri Jun 28 1996 09:288

CMP.B3  - Problems with wafer loss detection.  The tool was detecting lost
	  wafers while testing an engineering pad.  Attempted to jump out wafer 
	  detect sensors but this lead to invertor errors.  The tool was then 
	  returned to production w/ an IC 1000 installed and the sensors
	  adjusted accordingly.  Breakin wafers were run with no sensor errors.
     
21.32Table Gearbox Is Leaking AgainASDG::POIRIERSat Jun 29 1996 20:0110
    
    Oil Leak In Table GearBox:
    ---------------------------
    FYI: The oil puddle under the gearbox is continuing to grow. 
    
         It's now approx 6" in diameter.
    
                                               
                                                   C Shift.
    
21.33pad cond / scrub motorFABSIX::B_FINNWed Jul 03 1996 20:3614
	Scrub brush motor was intermittently not working. Checked 
cabled and brushes (motor), motor was shot. None in stock room 94 or 
cabinet. Kevin had one in F.S. office.  It was an upgraded motor like 
(B1 AND B2). Gear ratio is different. Installed motor and it screamed. 
Kevin came in to perform calibration to rpm. Procedure took longer 
than both of us expected. Aok now

Also checked Pad Conditioner down force. It was 89.9 PSI.  We adjusted 
back to 77 PSI.  Also asked manuf. tech to run MQC again. It passed 
fine.  Tim would like B1 and B2 Pad conditioner checked also as they 
become available. 12 lots in room now.

					A shift
21.34Start up B3LUDWIG::SWENSONFri Jul 05 1996 20:3310
    Jumpered di pressure switch
    powered system up with no problems
    no carriers were installed due to there were only 16 new carriers,
    cut primary pad.
    retrieved new brushes from stockroom for scrubber.
    
                                  C Shift
    
    

21.35Conditioner, downforce and scrubber issuesSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringThu Jul 11 1996 07:5941

       The downforce was checked on all (5) heads.  No problems
    found.  All heads were within tolerance except head #3 which
    out by 17 lbs. +/- 2 lbs.  Below is the chart results.  After
    the downforce was checked and MQC's were run, the tool was still
    in spec.  Next, I put the original conditioner wheel back on.
    The tool failed for fast removal.  Frank is being called as
    instructed.  We noticed scratches on the particle wafers.  While
    looking at the scrubber, it scrubs the wafer but stops during
    the cycle for (4) seconds or so before the scrubber drops down
    to come back.  It cycles fine with no wafer in it.  The 
    spacer configuration on the (4) posts that support the wafer
    after the pedestal places the wafer are not configured right
    but it looks like that it was put that way for a reason.  I
    took off the front bottom brush and the scrubber works fine.
    I left the bottom brush off and floating in the Elevator #2 
    platform for now until I speak with the Reps.

    HEAD          PRE                 POST
    ===========================================================
    #1       INPUT COEFF.--.47601    |  INPUT COEFF.--.45476
             OUTPUT COEFF.--5.87003  |  OUTPUT COEFF.--5.90981
             373 LBS                 |  376 +/- 1 LBS
    =========================================================== 
    #2       INPUT COEFF.--.47964    |  INPUT COEFF.--.45422
             OUTPUT COEFF.--5.81698  |  OUTPUT COEFF.--5.86737
             371 LBS                 |  376 +/- 1 LBS
    =========================================================== 
    #3       INPUT COEFF.--.50469    |  INPUT COEFF.--.50469
             OUTPUT COEFF.--5.83820  |  OUTPUT COEFF.--5.59947
             394 LBS                 |  376 LBS
    =========================================================== 
    #4       INPUT COEFF.--.47007    |  INPUT COEFF.--.46032
             OUTPUT COEFF.--5.77454  |  OUTPUT COEFF.--5.69231
             380 LBS                 |  376 +/- 1 LBS
    =========================================================== 
    #5       INPUT COEFF.--.46144    |  INPUT COEFF.--.46144
             OUTPUT COEFF.--6.03714  |  OUTPUT COEFF.--6.03714
             377 LBS                 |  377 LBS
    =========================================================== 
21.36B3 spin speeds checksSTRATA::KRUPAThu Jul 11 1996 18:2717
    
    Checked Spins speeds of table, conditioner wheel, and carriers. All
    appeared to match B2 except conditioner wheel.  Results were:
    					B3		B2
               		set-point	actual		actual
    	carrier		37		36.5		36.5
    	table		37		39		39
    	condition wheel	free		~29.5		~33
    Adjusted the postition of the conditioner from 9.0 to 10.0 and no
    change in speeds were observed.  
    
    Ran MQC's w/ no touchdown on 2nd table(like B1) and rate was
    ~3200A/min...compared to ~3400A/min w/2nd table.  2nd table may
    actually be removing a couple hundred A.  Need to further investigate
    oxide loss on the driven 2nd table.  Due to a lot of work in the room
    B3 was released to production until we have more time to run
    experiements.
21.37CARRIER#4 NOT ROTATINGSTRATA::WESTPHALThu Jul 18 1996 20:484
CMP.B3 - DOWN FOR THE CARRIER #4 NOT ROTATING.  THE INFRANOR WAS REPLACED AND 
	 THE TOOL WAS RETURNED TO PRODUCTION. 

21.38carrier one losing wafers/ down force issue.STRATA::SWENSONFri Jul 19 1996 20:0011
    1. carrier one not coming down for wafer clean.
    
    2. losing wafer out of carrier one.
    
    3. found the ep for carrier 1 to be misadjusted, adjusted the zero
       slightly which enabled carrier one to proceed down to wafer clean.
    
    4. The ep regulators need to be checked and down force recalibrated
    
     A call is being placed to speedfam for the analog output values and 
     corresponding pressures to correctly calibrate the ep regulators.
21.39Head #1 DF & Conditioner DF EP Regulators need Recalibration ASDG::POIRIERFri Jul 19 1996 20:5226
    
    Spindle #1 Down Force EP Regulator:
    ----------------------------------
    Upon completion of the Monthly PM, we noticed that Spindle #1 was not
    moving down to the scrub table with the other 4 heads. This condition
    led us to the zero pot on head #1's EP Regulator. A quarter turn on
    the zero pot made the head come down with the others. After further
    test runs we noticed head #1 repeatedly flying-out wafers at lift
    off. We attempted to calibrate the Ep regulator but could not find the 
    analog output values and corresponding pressure values needed to 
    properly perform the calibration.
    
    We called Speedfam and they agreed with our diagnosis. They are faxing
    us a calibration procedure in the morning.
    
    HEAD #1 HAS BEEN TAKEN OFF LINE UNTIL THE EP REGULATOR CAN BE
    CALIBRATED. PLEASE RUN WITH 4 HEADS UNTIL FURTHER NOTICE.   
    
                                                                   C Shift.
    Conditioner EP Regulator:
    ------------------------
    After completion of the Monthly PM we noticed the conditioner not moving
    down to the table on every cycle. We turned the zero pot on the 
    conditioner EP Regulator a quarter turn and the conditioner performed
    normally. This ep regulator will also be recalibrated tomorrow when 
    Speedfam faxes us the calibration values. 
21.40TROUBLE W/CARRIER#1 / SCRUBBER GONESTRATA::WESTPHALSat Jul 20 1996 20:399
CMP.B3 - Down the carrier#1 not going down at 2nd table and wafers sliding out 
	 on the primary table (same head).  Believe that the EP regulator is
	 out of calibration.  Waited for most of the day to receive a 
	 procedure from SF.  A procedure was found in the SF office and 
	 followed.  The down force for head #1 will need to be done over.  The
	 tool presently sets at PART waiting for the upper and lower scrub
	 assemblies that were put on CMP.B2.  Speedfam has been notified about
	 the scrubber and they are ordering the appropriate parts.

21.41water leakFABSIX::R_GEESun Jul 21 1996 08:097
	The system errored for a moisture sensor activated. 

	Founds a lot of water under the elevators. Cleaned the water 
	out repaired the leak under the #2 elevator and replaced the   
	moisture sensor cloth.

					B Shift
21.42DOWNFORCE AND EP REGULATORSTRATA::FINNSun Jul 21 1996 20:2458
	Repaired water leak at feed thru of elevator #1. Tested by dumping 
    filling and cycling elevator. Check throughout day, its leak tight as of  
    7pm.

    	Checked Pad condition downforce, based on C shifts comments. Found it 
    to be 90 PSI. Recalibrated, to 77 PSI.


	Checked EP regulator for Carrier #1, per procedure in F.S. office.
                                
    VALUE             ANALOG   CURRENT      PSI       DF

    0                        0        4        0        0
    500                   2020     5.95      6.2     53.6
    1000                  1828      7.9     13.8    142.5
    1500                  1636     9.85     21.3    243.5
    2000                  1445     11.8     28.8      333
    2500                  1254    13.76     36.2    421.4
    3000                  1063    15.71     43.6    510.3
    3500                   871    17.66     51.1    600.6
    4000                   679    19.62     58.5    690.7
    4095                   641    19.99    59.99      706

	Then performed downforce on Carrier #1. Found it low at 351. Adjusted 
    to 377. Then checked other carriers, and found them low, had to calibrated 
    all five  carriers.

	Carrier 1:	
			OLD					NEW

Input coefficient:	0.47903		Input coefficient:  0.45150	
Output coefficient:	6.79045		Output coefficient: 6.08753	

	Carrier 2:

Input coefficient:	0.48834		Input coefficient:  0.46601	
Output coefficient:	5.52785		Output coefficient: 5.79841	

	Carrier 3:

Input coefficient:	0.51432		Input coefficient:  0.47066	
Output coefficient:	5.59947		Output coefficient: 6.13793	

	Carrier 4:

Input coefficient:	0.47066		Input coefficient:  0.46716	
Output coefficient:	5.61008		Output coefficient: 5.66844	

	Carrier 5

Input coefficient:	0.48898		Input coefficient:  0.45476	
Output coefficient:	5.57029		Output coefficient: 5.98674	

Placed system in Wet idle.


					A shift
21.43SCRUBBER MOTORSUBPAC::CMPMon Jul 22 1996 16:108
    THIS IS FOR THE SCRUBBER PROBLEM.
    
    AFTER SPEAKING WITH THE TECHS OVER THE WEEKEND, ORDERED PARTS FOR A 
    SCRUBBER THAT WAS THOUGHT TO NEED REBUILDING.ON MONDAY FOUND THE MOTOR
    WAS THE ONLY THING THAT NEEDED REPLACEMENT AND THE SCREW THAT SECURES
    THE MOTOR SHAFT TO THE DRIVE ASSY.LOOSE TO THE POINT OF NOT TURNING
    THE BRUSHES. THERE IS A ??? OF WHETHER THE MOTOR NEEDED TO BE REPLACED
    OR JUST TIGHTENED TO THE DRIVE ASSY.
21.44Table spray nossle / replaced carrier #5FABSIX::R_GEETue Jul 23 1996 07:5010
	Replaced the primary and secondary pads. 
	
	Every run during the pad breakin a wafer was left on the 
	table by head #5 we replaced carrier #5. 
 
	Some how one of the hand spray nozzles used to clean the 
        table got caught by the table and ripped out. Had to repair 
    	the hose and replace all the table drain tubes. 
    
					B Shift
21.45SCRUBBER REPAIRSUBPAC::CMPThu Jul 25 1996 12:027
    REPLACED A FEW BEARINGS IN THE SCRUBBER THAT WERE VERY WORN, MADE
    BELT ADJUSTMENT FOR A VERY TIGHT BELT ON UPPER UNIT. I PUT B2'S
    SCUBBER BACK TO B2 AND B3' IS NOW RETURNED. I DID NOT NOTICE ANY
    UNUSUAL WEAR WITH COMPONENTS EXCEPT FOR THE BEARINGS THAT WERE 
    REPLACED.
    
                                        RICH (MR. GOODWRENCH)
21.46Warning status on scrubberFABSIX::L_SHKOLNIKOVFri Jul 26 1996 08:2011
    

    Scrubber wasn't working again this night.It scrubs the wafer but stops
    during the cycle before the scrubber drops down to come back. 
    It cycles fine with no wafer in it. I put warning on it. Looks like the
    lower unit going up too far. So I pulled out the scrubber assembly and
    found that the drive belt in lower unit was set too tight. I adjust the
    drive belt tension. cycle 25 wafers works fine so far. The tool passed
    MQC's and running product now.

                                                     Leon
21.47no slurry thru pump 2STRATA::RIDLONEliminate the obviousSun Jul 28 1996 05:3313
    Down for low removal rate from C-shift. They had replaced the pad with
    no change. We checked the carrier step heights and found 4 of the 5 to
    be off. Tightened the plate screws (1/8 to 1/4 turn) and the step
    heights came back in. Reran mqc's and the removal rate was still low.
    We then put on a precut pad and while the machine was running breakins
    the slurry looked to clear. Found that no slurry was flowing through
    pump 2. This was traced back to the incoming DI psi being to high at
    36psi which was not allowing the slurry pump 2 di rinse solenoid to
    close. Lowered the di psi to 30 and slurry flowed fine. We checked the
    low di alarm by dumping both elevators and the pressure didn't go below
    16psi. 
    
    					B-shift
21.48Mapper errorFABSIX::R_GEETue Jul 30 1996 07:525
		Came up with a mapper error, it wouldn't recognize
		cassette "A". powered down the mapper, powered
		up the mapper and homed the load module, all set.

					B Shift
21.49b3 scrubberSUBPAC::CMPThu Aug 01 1996 15:483
    checked and adjusted scrubber rpm
    
    
21.50unload flipper positioning and scrubber stationSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringFri Aug 02 1996 06:0211

           Two problems....

          1 - wafers being mis-handled at unload.  
              - retaught fingers at unload cup position

          2 - wafers cock-eyed at scrubber.  
              - one of the alignment spacers fell off it's
                post
                                         BJ
21.51PARTICLESSTRATA::WESTPHALSat Aug 03 1996 20:285
CMP.B3 - Down for high particles.  Changed both pads. Cleaned the conditioner 
	 ring and process area up.  Scrubbed the carriers. Returned the tool to
	 production.  It passed MQCs.

21.52Westec tie-in...SUBPAC::BJUBINVILLECMP-6 Equipment EngineeringMon Aug 05 1996 19:409
    

          Tool went down to EXTERNAL for facility tie-ins for
       CMP.C1-7.  Unfortunately, the GCW and other items are 
       tied into all of the tools in CMP-6.  A completed CMP 
       shut will be planned for the near future.  The weekly PM 
       was completed when the tool came back.

                                          A shift
21.53B3, sensors getting tired.FABSIX::L_SHKOLNIKOVWed Aug 07 1996 08:1119
    

         Tool alarmed numerous times for time-out waiting for
    scrubber to go to home, It wouldn't go all way back.
    Found scrubber home sensor (pushbutton) stack inside of the sensor 
    assembly. I removed sensor out of the tool took it apart and cleaned it.
    Sensor seal is broken and water getting inside. It is working now but 
    needs to be replaced.
    Later tonight we found some slurry in the slurry box slurry line #1 was 
    ready  to blow but the Moisture Sensor did not alarmed. We replaced the
    slurry lines and reset leak sensor module.
    Finally wafer detection system misalarmed twice for lost wafer from 
    carrier #5 so wafer detection sensor #5 has been adjusted - no problem 
    since than.
 

                                                      D - shift
    
21.54particlesLUDWIG::WESTPHALWed Aug 07 1996 20:458
CMP.B3 - Down for high particles.  The tool was at engineering most of the day.
	 Slurry was drawn directly out of a 55 gallon drum to eliminate the
	 slurry system.  The particles failed.  The pad was changed and MQCs 
	 rechecked.  The MQCs failed for particles again.  MQCs were checked 
	 using a recipe with an extended 2nd polish time but the tool still 
	 failed.  Please follow Engineering's action plan sent out by Krupa.

21.55PARTICLESSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringThu Aug 08 1996 07:5114


          C shift ran Frank's game plan...  still failed for
       particles.  New carriers were put on and a new pre-cut 
       pad from STK-6 was installed.  The 2nd table pad was 
       replaced as well.  The tool was thoroughly cleaned.   

       Break-ins are starting now.
                                   
       ps.. B2 passed paricles, lower than old spec values...

                                      D shift
                                             
21.56Table spinning out of controlSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringThu Aug 08 1996 08:0013



             When the OPS tech started running the tool for
         break-ins, the table started rotating fast while 
         the tool was loading up the first (5) wafers.  
         Leon was able to get out to DOS but the table kept
         spinning fast and wouldn't stop.  The tool was then
         softbooted and it kept running.  The tool ended up
         being EMO'ed.  It came up fine after that.

                                             D shift
21.57PARTICLESSTRATA::WESTPHALFri Aug 09 1996 20:2214
CMP.B3 - Down for high particles and testing by process engineering.  For the 
	 full passdown please refer to Tim Dyer's passdown.  Had some handling 
	 problems with the unload flipper manually exercised the flipper and no
	 further problems were experienced.  Started to follow engineering's 
	 particle get well plan.  Ran the test using the unfiltered fab4 
	 slurry.  The results were as follows:

                                PRE              POST
                          ---------------------------------
                          1. 13-37-0      |   343-1344-0
                          2. 11-35-0      |   272- 918-0
                          3.  6-34-0      |   333-1303-0


21.60B3 crashFABSIX::J_SWENSONSat Aug 10 1996 20:378
    Tool crashed 30 seconds into segment 6.
    replaced carriers 1,2,and 5.
    cleaned tool and reset counters.
    returned system back to porduction for Mqc's.
    
    
    
                               c-shift
21.58Door rollersFABSIX::R_GEETue Aug 13 1996 08:2111
	The door interlocks kept stopping the system. Found all
	4 process doors were off the tracks. Most of the screw holes 
	that hold the door rollers on were stripped. Tapped out the hole
	on the top of the door to 10-24 and reinstalled the rollers.
	We did this only to some of the top holes, didn't want to due
	this to the bottom rollers because they won't roll only slide.

	This is only a short term fix, we need new door rails to repair 
	them correctly.

					B Shift
21.59belt skipping.STRATA::RIDLONEliminate the obviousWed Aug 14 1996 07:507
    Polish table belt slipping on bkin3 recipe. Slurry was suspect. Ashland
    was changing the filter at 6am and the last time the filter was checked
    was at 8pm. However after the filter was replaced 5 more breakins were
    run and the belt still slipped, only once though. Leaving in wet mode
    for A-shift eval. 
    
    						B-shift
21.60loose beltSUBPAC::CMPThu Aug 15 1996 14:433
    Found that the belt from the motor to the center shaft was loose,(the
    bolts that lock the hub in place were loose) adjusted the belt tension
    and retightened the hub bolts.
21.61loose beltSUBPAC::CMPThu Aug 15 1996 14:462
    Found that the belt from the motor to the center shaft was loose( bolts
    on the hub) readjusted the tension and retightened the bolts.
21.62Vacuum pump movedFABSIX::R_GEEMon Aug 19 1996 08:347
    Relocated vacuum pump to the black box below the floor.
    Test fine.
    
    
    					B-shift
    
    
21.63a multitude of problemsSUBPAC::LANDRYMon Aug 19 1996 20:1327
    Several problems with this system today............
    
    *  The Unload Flipper was alarming for "no wafer present", even though  
       it possessed a wafer.  The alarm was happening on every wafer, very
       consistant.  Checked the I/O (with a wafer being gripped) for proper
       functionality.  The LED remained off.  Found that the travel sweep
       of the gripper to wafer was excessive.  Observed that the I/O (#7)
       would flash as the grippers were closing but once gripped, the I/O 
       was de-luminated.  Therefore, made a mechanical adjustment of the 
       grippers to reduce the distance of the grip sweep.  Cycled many
       times, proper I/O sequencing each time.  Also adjusted the flipper
       to the unload cup position to better grip the wafer as it was
       raising the wafer when gripping.
    
    *  The screen locked up today while trying to unload wafers.  This
       occurred when the above problem was happening.  Had to reset the CPU
       to correct the problem.
    
    *  The system later on alarmed for "low DI water pressure".  Checked
       the incoming supply and the internal gauge.  Both were reading fine.
       Hit a key to continue and the alarm never re-appeared.  BUT, the Pad
       Conditioning sequence remained on (virtually locked up).  We were
       able to exit the "run" screen, go into manual, and end the pad con-
       ditioning segment.  Went back into the "run" screen and restarted.
       No problems since that.
                                            A shift
    
21.64annoying little things, Part IISUBPAC::LANDRYTue Aug 20 1996 20:3126
    More annoying little problems today........
    
    *  The system powered down today for no apparrent reason.  Had the 240
       VAC coming into the machine, but unable to power up the system.  
       Traced the problem to the Inverter Cabinet interlock.  Repaired the
       problem and power up was accomplished.
    
    *  From there, the system was getting "process door open", not able to
       initialize.  Found several switches not making it due to poor door
       rollers.  Fixed those and away we went.
    
    
    *  Next up was a couple of second table carrier rinse valves staying on
       when they should have shut off.  DI water everywhere!!!  Even out
       the top of the system onto the floor.  Found the incoming pressure 
       up over 30 PSI.  This high pressure is too much for the valves to
       handle.  Reduced the pressure to 24 PSI for the fix.
    
    *  Again today, the DI water alarm went off for no apparrent reason. 
       As was the case yesterday, it happened when the system was in a Pad
       conditioning cycle.  Hit a button to continue and off we went. 
       Again like yesterday, the conditioning cycle locked up, requiring us
       to exit out of the screen to stop the cycle.  
    
                                        Doc
    
21.65Wafer Loss Detection Sensors AdjustedASDG::POIRIERFri Aug 23 1996 20:1711
    
    Wafer Loss Detection Sensors:
    ----------------------------
    The wafer loss detection sensors were adjusted due to numerous
    assists for wafer loss errors. 
    
    
    
    
                                                  C Shift.
    
21.66MAPPERS GROUNDEDSTRATA::RIDLONEliminate the obviousTue Aug 27 1996 06:367
    Mapper intermittently failing while trying to map cassettes. Powering
    down and back would correct. We installed ground wires to both mappers
    to eliminate any possible static build up problems. Hasn't failed since
    but it wasn't happening all the time to begin with.
    
    
    						B-shift
21.67Mapper Timeout / Scrubber Timeout Errors!ASDG::POIRIERThu Aug 29 1996 20:2726
    
    Mapper Error:
    ------------
    The system errored for "Mapper Timeout" this morning. The Cassette 
    mounting base would tilt and relevel itself properly but the mapper
    would only rise to the first slot then stop. After 4 repetitions, the
    mapper would timeout. The program was exited to DOS, the Kensington
    Controller was reset and a soft boot was performed. The mapper operated
    normally for the rest of the shift.
    
    The Speedfam rep ordered a replacement Kensington Controller as a 
    precautionary measure.  It is expected to arrive tomorrow 8/30.
    Based on past performance history, Richie believes the Controller  
    is apt to fail soon.
    
    
    Scrubber Movement Trouble:
    -------------------------
    The system was erroring this afternoon for "Scrubber Timeout". The
    scrubber was hanging up during its in/out travel. The rails were 
    cleaned & lubricated but the scrubber still did not move as smoothly
    as we wanted. The flow controllers were opened up slightly and the 
    motion returned to normal.
    
                                                              
                                                          C Shift. 
21.68Scrubber Timeout Errors / Dampners Removed ASDG::POIRIERFri Aug 30 1996 20:3813
    
    * Scrubber Timeout Problems:
      _________________________
      The Scrubber Dampners (Air Shocks) were removed from Scrubber today.
      The Speedfam rep tells me that they were removed from B1 & B2 as well.
      It appears that they are more trouble than they're worth. The front
      dampner had lost its ability to absorb shock and was causing the
      scrubber to bounce back off the "scrubber in" sensor.
    
      Both front & rear shocks were removed and the flow controllers were
      adjusted for slow movement to prevent slamming at the end of the 
      stroke.
    
21.69Load Flipper Failing To Pick Up At Pre AlignerASDG::POIRIERSat Aug 31 1996 20:319
    
    Load Flipper Problems:
    ---------------------
    The Load Flipper failed to grasp a wafer from the pre aligner 3 times
    today. We weren't able to observe this condition, and since the tool was
    run for most of the shift, we didn't get an opportunity to investigate. 
    
    
                                                                C Shift.
21.70INVERTER ERRORSUBPAC::CMPThu Sep 05 1996 17:392
    TOOL HAD AN INVERTER ERROR, RESEATED THE CABLES FROM INVERTER TO
    COMPUTER AND BROUGHT BACK UP.
21.71Inverter Communication ErrorsASDG::POIRIERThu Sep 05 1996 20:229
    
    Inverter Timeout Errors:
    ------------------------
    Inverter Timeout errors repeatedly occured today during Wet Mode.
    The series #10 computer cable next to the Inverter was snugged.
    The system has been in Wet Idle since 3:00 with no further errors.
    
                                                         C Shift.
    
21.72Mapper ErrorFABSIX::R_GEETue Sep 10 1996 08:395
	The system developed a Mapper error, tried several times
	to map the cassettes with no luck. Powered down the mapper
	and re-homed the load module, all set for now.

					B Shift
21.73Robot to cassette adjustmentFABSIX::R_GEEWed Sep 11 1996 08:287
	The robot was intermittently missing wafers at cassette #1,
	found the robot to cassette alignment was off.

	We adjusted the robot position and cassette tilt for both
	#1 & #2 cassettes.
 
					B Shift
21.74Carriers changed/System lockupLUDWIG::SWENSONFri Sep 13 1996 10:144
    Changed carriers 1,2 and 3 due to high wafer count.
    While in service mode tool locked up for timeout for process door 
    inter-lock. Tool needed soft reboot to rectify lockup.
    Returned tool back to production.
21.75Carrier clean brush movedSTRATA::WESTPHALFri Sep 13 1996 19:504
    CMP.B3 - Moved the carrier clean brush about an inch counter clockwise
    	     as requested by engieering.  This was done to increase the
    	     brush to carrier contact.
    
21.76switch fixSUBPAC::CMPWed Sep 18 1996 19:365
    Had an inverter error today, the possible cause could be the darn doors
    not allowing the interlock switch to close, so when in wet mode and the
    pad arm is ready to drop on the polish table, the table does spin
    causing the error. will look into enlarging the platform for the
    switches.
21.77Inverter error during wet idleLUDWIG::SWENSONWed Sep 18 1996 20:118
     To follow on last reply doors will be reworked also.
     Also, to bring the tool up from the inverter errror
     we exited to dos, powered tool down and inspected interlock switches.
     found no switches to be out of whack. powered tool up, no problems.
     returned system to production with a warning of the inverter error.
    
                          
                           C-shift
21.78Problem during PMFABSIX::O_WHITFORDThu Sep 19 1996 07:586
         During the "look see" phase of the weekly PM, the moisture alarm
    went off. We discovered a lot of slurry leaking from pump #1. We
    replace hoses for pumps 1, 2 & 3 and checked flow rates which were in
    spec.
    
    Dave & Leon
21.79INVERTER PROBLEMSFABSIX::J_SWENSONThu Sep 19 1996 20:1928
          Inverter error first occurred this morning. Tried troubleshooting
          what could be and tool ended up being powered down and up and
          error cleared as usual. Check 3 phase in and out of the inverter
          no problems. Checked all door interlocks and made some enhancements
          but none were tripped.  Checked and found the connection from RS485
          connecter to the RS232 connection to be very loose at the side of the
          cabinet (no posts between connectors). Installed mini tester between
          the two and and secured. (observed transmit signal but no recieve 
          signal, the tool was working properly  though. Completed weekly
          p.m. after seeing there was no more visible problems and the tool
          had been rebooted. Returned system to manufacturing, mqc's were run
          without incident. The  first five monitor breakin's ran fine and upon
          the next set of five the inverter error orcurred in segment 3 of
          recipe ILD1, table was not spinning, the carriers were spinning and 
          being lowered to the table so tool was emoed and rebootted. During 
          reboot noticed we did not have inverter microprocessor or the +15 
          volt from power supply number six. Reseated inverter cover and 
          regained micro processor and when measuring ps 6 again  +15 was there.
          This cleared the error again still no solution. Removed RS232 tester
          secured connectors in PD cabinet. Checked RS232 connection at xport
          number seven traced ribbon cable to scm mod no visual problems.
          Left pd cabinet open with cone in front for future trouble shooting.
          Ran 25 wafers with same program ILD1 without incident and returned 
          system to production. 

                                 c-shift,Kevin and rich
 
              
21.80invert comm errorsSTRATA::RIDLONEliminate the obviousTue Sep 24 1996 05:5710
    Operator was supposed to put tool down for load cell error. Screen
    locked up and table was spinning. Had to emo tool to recover. Load cell
    problem could not be duplicated. But in investigating that, we ended up
    having continuous inverter comm errors. Opened up the inverter and
    found RT1 broken. RT1 in on the RS232 line. Ran 15 wafers using
    bkin2, bkin3, and bkin. Time will tell if this is the fix or not.
    
    
    						B-shift
    
21.81vacuum problem resolvedSUBPAC::LANDRYTue Sep 24 1996 20:2924
    Continuation of the previous shift recovery......
    
    The system was alarming for "Low Main Vacuum".  When in idle or wet
    mode, the gauge off the pump was at ~27" Hg, excellent.  Tried running
    the NO_TOUCH program and observed the I/O's for carrier vacuum.  It was
    noted that head five was the first carrier to alarm, with the remaining
    four heads not far behind.  So, we dis-abled head five and ran the same
    program with the other four heads.  No error.  Dropped carrier five and
    inspected the carrier.  The were some major smudges and "gritting
    particles on the film surface.  Installed a new carrier and retested
    several runs.  Good results across all five heads.
    
    Whilst the carrier was being removed, the heads were flushed out and
    the vacuum tanked dumped several times to assure clear sailing. 
    Breakin and MQC's went well.  No errors were observed during running
    today.
    
    After the system came up, the carriers from the previuos night were
    dismantled and checked.  Major, no MAJOR, black pad mulch everywhere
    in the carriers.  Cleaned out all affected components and re-assembled
    the carrier bodies.   
    
    					Doc  /  Kevin C.
    
21.82Carrier crash on 2nd table 9/23FABSIX::R_POIRIERTue Sep 24 1996 23:4314
	Operations was running a lot and paused the tool to replace a
	wafer in the index table. When they resumed, the Multihead came
	down on the 2nd table with no wafers on the carriers, started 
	rotation when carrier 3 stopped.
	
	When carrier 3 stopped rotating it tore up the pad. All carriers, 
	and both pads were contaminated with pieces of the black pad. 
	We cleaned both tables and replaced both pads and all 5 carriers. 

	
    
    						B Shift
    
21.83Pre Aligner Sensor Failure / Tin Split LotASDG::POIRIERThu Sep 26 1996 19:4210
    
    Pre Aligner Sensor:
    ------------------
    
    The sensor was failing to detect wafers during a (Tin) split lot run.
    It appears that the reflectivity differerence of the Tin was enough
    to effect the sensor. We moved the sensor closer to the wafer surface
    and had no further trouble. 
    
                                                           C Shift.
21.84groundingSUBPAC::CMPFri Sep 27 1996 17:091
    The grounding of the robot and mapper is now completed.
21.85Carrier Rinse For Head #2 Failed to Turn Off.ASDG::POIRIERSat Sep 28 1996 20:2512
    
    Wafer Loss Detection Error:
    --------------------------
    The system errored for "Wafer Loss" today. The Carrier rinse water
    nozzle for head #2 failed to turn off. Water was sprayed on the 
    wafer loss sensor between heads 1 & 2 causing the error.
    Keep an eye on SV62, we may want to replace it if this happens again.
    
                                                                 C Shift.
    
    
      
21.86I thought the bugs were gone.LUDWIG::RIDLONEliminate the obliviousSun Sep 29 1996 23:0016
    System sitting there in segment 12 of tup_f after shift change. 2 wafer
    had loaded, 1 in load flipper, 1 in aligner, and 1 on robot. 4 wafers
    in unload cups with 1 unloaded. 5 wafers on carriers that appeared to
    have polished. Noticed the index table wasn't where it should be. It
    didn't rotate far enough so that the unload and load cups were under
    the solenoid. Not sure if this was the cause of it stopping or the
    result of another problem as the machine wasn't alarming at all. Tried
    to exit out and restart but maching wouldn't count down. Tried homing
    the index table and indexing to proper cup but still a nogo. Tried
    homing the mhta but still nothing. Checked the Digins and those should
    have been lit were and those that shouldn't have been lit weren't.
    Instructed operator to load up 10 wafers, running the first with tup_e
    and the second with tup_f (the to recipes involved with this fiasco).
    
    
    						B-shift
21.87what the heck?LUDWIG::RIDLONEliminate the obliviousSun Sep 29 1996 23:258
    The 10 wafers ran with no problems. The room is short on bkins so we're
    going to leave this at that and shrug the "fault" off as a lost bit.
    NOTE: When viewing the errors.log file under the view ascii file button
    you only see faults up to July 12th (or there abouts). What the heck?
    Doesn't offer much help when trying to troubleshoot.
    
    	
    						B-shift
21.88#1 carrier clean nozzleLUDWIG::RIDLONEliminate the obliviousMon Sep 30 1996 05:597
    Two wafer loss sense errors while processing. Both times it was due to
    #2 wafer loss sense being wet. Dried and resumed both times with no
    problems. ONce lot was done repositioned #1 carrier clean DI nozzle
    which wasn't centered on carrier and overspraying. 
    
    
    						B-shift
21.89inverter errorSUBPAC::CMPTue Oct 01 1996 20:517
    Had an inverter error today and investigated. We found we had lost the
    +15v on the power supply so we took the leads off TB9 and found the
    voltage returned so we put the wires back only to find the voltage
    was still there. We checked all the +15v connections at the TB's for
    tightness and are still monitoring, with no error yet. I would suggest
    that if this happens again that the +15v gets checked and look for
    something that may be dragging the voltage down.
21.90Mapper ErrorsFABSIX::J_SWENSONSat Oct 05 1996 20:317
      Had a mapper error and could not clear.
      checked all associated cables and grounds.
      found exposed ground wire and fuse chain on back of controller
      to be touching. taped up ground and monitored mapper error,
      Had to soft boot to clear error.
     no error's  ocurred rest of shift. assist was logged.
                
21.91pad changeLUDWIG::RIDLONEliminate the obliviousWed Oct 09 1996 08:0311
    A funny thing happened on the way to the unload cups....it dropped the
    wafers. The subsequent main vac alarm was no surprise but when the
    wafer track said the cups had unloaded AND then reloaded that was,
    seeins how the load cups were still full. Anyhoot, we loaded the
    carriers in manual ops and then restarted the 1 second touch ups and
    observed while the system polished the next 10 wafers with out a hitch.
    After the 15 wafers were completed it was time for the scheduled pad
    change, 2001 pri and 1991 sec. Returned to production for mqc's.
    
    
    						B-shift 
21.92Low DI Water Pressure errorPATE::GEEMon Oct 14 1996 20:058
	The system errored for "Low DI water pressure". When we tried
	to recover the touch screen froze-up on use, had to reset the
	computer to recover. After the system recovered we checked
	the DI water pressure at the systerm, it was 18 lbs. Adjusted
	the main DI water regulator up to 24 lbs and ran 2 verification
	runs, all set.

				A shift
21.9315v Power supplyFABSIX::L_SHKOLNIKOVThu Oct 17 1996 06:429
      
   The system was alarming for inverter error. Operator EMOed the tool.
   After we powered it up I checked the 15 v power supply and did not 
   get +15v. After the output leads were disconnected the voltage returned.
   Put the wires back - 0. So 15 volts power supply has been replaced.
   System came back up and running MQC's now. No errors so far.


                                                       Leon
21.94Mapper Errors - Please Document All FailuresASDG::POIRIERFri Oct 18 1996 20:2812
    
    * Mapper Errors:
      -------------
      The system errored for the mapper this morning. The system was soft 
      booted and the kensington mapper controller was powered down. The
      mapper performed normally for the rest of the shift. This is the
      3rd time this week (on C Shift) that the mapper has failed. Please
      make sure to log all mapper failures on CMP.B3 in Notes. We need the 
      documentation to convince Speedfam to replace the controller. The
      Speedfam reps have a spare controller in their office.
    
                                                               C Shift. 
21.95ionizer installedLUDWIG::RIDLONEliminate the obliviousSun Oct 20 1996 08:128
    John Swenson passed on that he had obtained a ionizer for this tool but
    was unable to install. Apparently some other companies have experienced
    mapper errors as we have and an ionizer remedied the problem. We were
    able to install as the tool wasn't used to capacity. Please continue to
    document any mapper errors.
    
    						B-shift
    
21.96vac/particle problemsFABSIX::R_POIRIERMon Oct 21 1996 05:3911
    Failed for particles and also had some vacuum problems. We found the
    wafer in carrier 2 misaligned but also carrier 3 vacuum switch would
    not make. Cup to head alignment was checked and adjusted accordingly.
    Carrier 3 was swapped with carrier 5 and during no touch carrier 5's
    vacuum switch didn't make. Carrier 5 was changed and mqc's ran which
    failed for hi particles. Inspected the pads and found red particles on
    the primary platen. Changed primary and secondary pads. Over to
    production for bkins and mqc's.
    
    
    						B-shift 
21.97particlesSTRATA::RIDLONEliminate the obliviousTue Oct 22 1996 06:1511
    AFter Ashlands work on the slurry system this tool failed for particles
    X2 again. Point of use filter (1 micron) was installed and slurry flow
    checked. During mqc's it crashed while running breakins. Both lines on
    the pumps blew. Apparently the 1 micron couldn't handle the flow and
    pressure built up at the pumps until the lines burst. The pad and three
    carriers were changed as well as the pump lines. The pads been
    conditioned and awaits further evaluation from EE. 
    
    
    					B-shift
    
21.98Down force cal.FABSIX::R_GEETue Oct 22 1996 20:0312
	Work done on 10/21.

	The system failed MQC's multiple times for high particles and
	also head to head removal was out. 

	Per Eng. we calibrated down force on all 5 heads, all heads needed 
    	some adjustments. Replaced the primary and secondary pads, performed 
	the weekly p.m. and cleaned the system thoroughly. 

                  				A shift

						A shift
21.99leakSTRATA::RIDLONEliminate the obliviousTue Oct 22 1996 23:0411
    Failed uniformity X2 on heads 3 and 4. While changing them out we
    noticed that the slurry pumps were soaked. The leak detector was also
    soaked but still green (indicating no leak). Replaced the hose for pump
    1 (split). Both hoses were just replaced yesterday so pump 2 was not
    changed. The leak detector had a piece of nylon wipe under it that was
    soaked. We tested the sensor with the standard paper (that supposed to
    be there) that was soaked and it worked fine. Tested pumps after and
    they worked fine and dry.
    
    
    					B-shift 
21.100jump on itSTRATA::RIDLONEliminate the obliviousWed Oct 23 1996 07:5816
    During breakins for mqc's wafers we got a wafers not on carrier alarm.
    Found wafer on head 2 not seated. Reseated and continued. Next set of 5
    crashed. A wafer came out of 2 and took out 1. Replaced the 2 bad
    carriers and the pad. On the subsequent mqc tests the tool failed for
    particles. All wafer showing the tell tale signs of slurry build up
    around the edge. REplaced the brushes after the first test due to them
    looking pretty bad. The second mqc didn't fair much better. Fearing we
    have a major problem with the slurry delivery lines a barrel was
    brought up and hooked up locally. Flowed slurry to test and it worked
    fine. Now just waiting on ops to measure up some more mqc wafers. Other
    tools didn't fail because they weren't tested. This will continue over
    to A-shift. 
    
    
    
    						B-shift
21.101cup 2STRATA::RIDLONEliminate the obliviousWed Oct 23 1996 08:0410
    Forgot to mention that we rechecked the carrier to cup alignment and
    cup height alignment. Both were good. We slowed down the down motion
    for cup 2. We also checked the solenoid/plunger for any obvious signs
    of evil there. Found that the plunger was not level as was evident by
    the marks on the bottom of the plunger (as well as the level we used).
    Rotated the plunger about 225 degrees and it leveled itself. Didn't
    have any more loading problems the rest of the night.
    
    				
    						B-shift
21.102Slurry pump 1 line brokeFABSIX::R_POIRIERWed Oct 23 1996 08:119
    Peristalic pump 1 line cracked and was leaking. It didn't alarm because
    of the material I used for a pad under the detector. We replaced the line 
    and the pad and cleaned the area.
     
    Later in the shift line 1 broke again causing an alarm. Again we replaced 
    the line and the pad and cleaned the area.
    
    
    						B Shift
21.103Slurry & DI clean linesFABSIX::R_GEEWed Oct 23 1996 20:3316
    
	The system failed MQC's for high particles again. We calibrated
	the slurry flow for both pump #1 & #2 individually, both were out
	slightly. We tried to then due a calibration check with both pumps
	running (551 mlm Standard process), it was way bellow the requested
	flow rate, we found this was also the case with the table rinse
	since they use the same lines.. 
	We found that pump #1 & #2, 1/4" lines were tee'd together after 
	the pump, this wouldn't allow the correct flow to pass at the
	standard hi volume. We removed the tee and connected the unused
	pump #3 line to pump #2. 
	Now pump #1 & #2 have there own individual slurry dispense lines.
	Calibrated the slurry flow and everything looks good. The unused
	Pump #3, 1/4" line has been plugged after the pump.

						A shift
21.104Recipe issuesSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringThu Oct 24 1996 08:1112
    
    
         Tool kept leaving wafers at lift off as a result of
       the new recipe.  Vacuum and carriers were verified as
       not being the problem.  Spoke with Frank and had us go
       back to edge lift off which he said will probably show
       particles, which it failed twice for.  I tried extended
       rinse times and speeds but I couldn't get the center
       lift off to come through.  It 7am now so I'm going to
       wait for ENG to come in.
    
                                                 BJ
21.105particles/double slotted wafers/lift off problemsFABSIX::J_SWENSONThu Oct 24 1996 19:3816
    1. System went down for slurry pump 3 hose (replaced)
    
    2. system intermittently  double slotting wafers at recieve cassette #2
       ( ongoing )
    
    3. Changed out slurry supply tubing to muultihead (slurry flow
       calibrated)
    
    4. changed primary and secondary pads @ 1800 per Frank Krupa.
    
    3. Process engineering continuing to evaluate wafer droppage/particles
       during lift off.
    
    
                                  C-Shift
      
21.106B3 tail chasing updateSTRATA::WOOLDRIDGEPleasure, Spiked With PainThu Oct 24 1996 20:1472
      B3 particle problem update 10/24:
!*see end for plans & status:

Teflon liner tubing mqc:

std recipe/ segment 6, 10s-377, seg 7, 20s 313, lift 12.7 

Particle results:
large     tot   scratch
1 - 274   989   5.10
2 - 294   2286  32000
3 - 342   1307  1.18
4 - 466   1727  2.32
5 - 479   2069  29.1

Same recipe, but center lift:

Particle results:
large     tot   scratch
1 - 446  1641  0
2 - 302  1258  0
3 - 522  1988  1.04
4 - 707  2525  1.23
5 - 651  2472  0


changed pump 3's tygon tube because we noticed it was leaking/cracked.  
Ran edge lift again-same recipe as above.

Particle results:
large     tot   scratch
2 - 481  1944  1.09
3 - 375  1138  0
4 - 264  8882  19.1
5 - 576  2228  4.13

Replaced teflon lined tube with old style

Particle results:
large     tot   scratch
1 - 227  724  1.24
2 - 297  1281 16.0
3 - 512  1929 0
4 - 244  901  6.94
5 - 194  674  26.3

Changed pads- cleaned conditioner- center lift off- same recipe:

Particle results:
large     tot   scratch
1 - failed 
2 - "
3 - "
4 - "
5 - "

----------------^^mqc's on B2 now have failed.  Although the 
Strasbaugh tools MQC's are still clean, the only way we now see 
getting the Speedfams up is to shut the room down and purge the slurry 
lines. Last time this was done, the lines were purged for an hour.  We 
are contacting Ashland now to begin purging the lines from 8pm to 
11pm - 3hrs.  After this, the tools should be MQC'd and we will see 
where we are at.

B3 qual should use the IO_Mqc recipe first. If this fails, try the 
CEN_MQC recipe. 

* Note- The scrubber on B3 is not working and needs repair.
        Try to repair this when the slurry system is being purged.

   -= Kyle & Frank
21.107Scrubber not workingSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringFri Oct 25 1996 05:009

            a report that the scrubber wasn't working.
         Found that the process time on the SCRUBBER #1
         recipe was set to 0.  Set to 7 seconds like B2.

                         and that's all she wrote!

                                             D shift
21.108POOR HEAD TO HEAD REMOVALLUDWIG::WESTPHALSat Oct 26 1996 20:4315
CMP.B3 - Down for head to head removal being out (head 2).  Swapped the carrier
	 and process rechecked the MQCs.  They failed.  The DF was checked.  
	 The EP regulators and DF were adjusted for each head.  

                            PRE        POST
                            381.5      377.2           
                            371.9      377.2
                            373.1      377.8
                            369.9      376.9
                            371.2      377.0

         The tool was set to verify-U to have the MQCs rechecked.  The MQCs 
	 passed so the repair was ended.

21.109crashLUDWIG::RIDLONEliminate the obliviousSun Oct 27 1996 06:2213
    Crashed during tup-175. 1 second into segment 5. Wafer came out of 2
    and took out 1. Alarm on screen was main vac low. Operator put that the
    alarm was wafer loss detect on the crash sheet but thats not the alarm
    Roger or I saw. Pad was changed and the two carriers. Checking cup to
    carrier alignments now. We are assuming that it was misload at the
    carrier due to the same thing occurring during our shift last week. Why
    it didn't happen on the other shifts is unknown. On a side note we
    observed the scrubber not operating properly. The brushes stopped
    prematurely while manually unloading 5 wafers. Motor was checked and
    smells like it could be bad. Maxim is down so this might be passed on. 
    
    
    						B-shift
21.110Index table alingmentFABSIX::R_POIRIERSun Oct 27 1996 07:279
    We found head 2 to be out of alignment. Also, Index table needed
    alignment. Old reading for Index table was -187. New reading after
    alignment is -288. 
    
    Conditioned pad and wet down carriers and returned tool to Operations.
    
    
    						B Shift
    
21.111Particles....resolved?SUBPAC::BJUBINVILLECMP-6 Equipment EngineeringWed Oct 30 1996 07:2520

         The "plan" went through all bad.  The last action on
      list is what was left which we ran NOTOUCH which looked fine.
      So, we went with an external batch of slurry from the barrel
      in the B3 chase.  It was lot # 12-17-96-22.  The tool passed
      with all heads under 50 particles.  The house system was 
      reconnected and it failed bad, 1500-2000.  The A tools passed
      particles fine as well as B1.  B3 is at the end of the loop so
      it could be just that, last in line gets the crap.  The work
      load is too great right now so OPS won't let us have a flush
      done.  The only thing we did notice is that the 2nd table pad 
      was changed and not the primary.  Per process spec, the 
      primary pad was change.  I don't expect to see anything good
      come out of this but it was just item that was open.  Break-in
      and MQC particles wafers were run and it passed particles with
      the rest of the MQC's running now.  Go figure...


                                             D shift
21.112Scrubber issueSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringThu Oct 31 1996 04:5311


              The scrubber stalling...  I was watching it and
          it does stop towards the end of the scrub process 
          and everytime it does, the MHTA has the heads on the
          table rotating.  When the heads are up and NOT rotating,
          the scrubber works fine?????????????????????
                                             

                                             BJ
21.113removing wellLUDWIG::RIDLONEliminate the obliviousSun Nov 03 1996 05:0411
    Low removal rate when we came in. C-shift was in the process of
    changing the pad so we finished it up. During the change we noticed
    that the carrier clean nozzle for head 2 was spraying. Found the di psi
    in the liquid cabinet to be at 26psi. Adjusted it down to 24 and
    spraying stopped. Checked the slurry cal and also checked the ph and
    both were good at 500mls/min and ~10 respectively. Removal rates came
    in on the first test run, uniformity on the third.
    
    
    						Combined C and B-shift
    
21.114late entry, flipper tubbingSTRATA::TDYERThu Nov 07 1996 15:4512
CMP.B3 - Down for particles and problems with the load flipper.  A barrel of 
	 slurry was hooked up at the tool with no recirculation loop installed
	 and the particles were rechecked.  The particles came in in the 100s.  
         Process Engineering wanted the MQCs rechecked with a recirculation
	 loop installed.  While check the MQCs it was noticed that the conduit
	 on the load flipper was broke leaving the wires exposed.  The conduit 
	 was cut back and reattached.  There was the later a problem with the 
	 positioning of the load flipper to prealigner.  This was probable 
	 caused when working on the conduit.  The flipper alignments were
	 performed and checked.  The tool sets needing only MQCs.

21.115Polish table gear box getting worseSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringFri Nov 08 1996 06:5811



            Whilest the weekly PM was being completed, Leon
         discovered that the surface for the drip-pan under
         the main polish table was almosty entirely covered
         with oil.  E.E. will be addressed in the morning 
         to check on status of tool.  A WARNING was placed on
         the tool.
                                             D shift
21.116Scrubber WARNINGSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringSat Nov 09 1996 06:368

            The scrubber is intermittently stopping during the scrub
        process again.  This is an on-going issue.  Due to the workload,
        we can't rip this puppy apart.  As soon as the B2 or the workload
        lightens up, it needs TLC.  I put a WARNING on the tool.

                                             D shift
21.117Mapper errorFABSIX::B_FINNMon Nov 11 1996 19:085
	Mapper error on B3. Ionizer turned back on. Fix has been 
        confirmed. Also needed a soft boot.

				A shift
21.118Pad changeSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringFri Nov 15 1996 06:226

          Tool failed MQC's for range and deviation.  Both
      pads and all (5) carriers were replaced.

                                             D shift
21.119Pad change due to bad unif.SUBPAC::BJUBINVILLECMP-6 Equipment EngineeringThu Nov 21 1996 05:437

         Tool uniformity went south during product lot.  We
      took the tool for pad changes and weekly PM.


                                             D shift
21.120READJUST SCRUBBER AMPSUBPAC::CMPThu Nov 21 1996 17:144
    Looked at the scrubber for a delay in the time the brushes were
    spinning. Found that the........... AMPLIFIER CARD WAS COMPLETELY
    MISADJUSTED...................readjusted and time was set for 11
    seconds. Tested AOK.
21.121Particle Control effortsFABSIX::O_WHITFORDQuticherwhinin&#039;Fri Nov 22 1996 21:1310
    Previous shift had high particles. Secondary pad had been changed and
    some minor mechanical tweaking to the DI delivery on the secondary
    table, all of which helped, but didn't completely control the
    particles. We replaced the primary polish pad and scrubbed down the
    polish table area including the pad conditioner ring & arm. Also
    drained the index table, scrubed the load and unload cups and wiped down
    the interior surfaces of the index table. Returned to ops for
    verification/MQC's.
    
                    D shift 
21.122Particles...SUBPAC::BJUBINVILLECMP-6 Equipment EngineeringSat Nov 23 1996 05:2212
            Tool kept failing for particles (large out of spec).
       The pads were replaced, tool cleaned ect...  The brushes
       were inspected.  NOTOUCH was run without any problems.
       Slurry flow was checked and it was ok.  Rinse was checked
       and that was fine.  An EXTERNAL barrel was hooked up and 
       particles passed.  The house system was put back on and
       it failed for large particles and over-all count.  We
       are awaiting for an E.E. call-back to ok a slurry flush
       from Ashland.
      
                                             D shift
21.123ParticlesFABSIX::J_SWENSONSat Nov 23 1996 18:4712
    
     Tool failed for particles, carried over from D-shift
    
     PAL filter changed over to Osmonic's
    
     One hour time before next mqc's
    
     mqc's passed for particles
    
     Tool released for production use
    
                         C-Shift
21.124Pump #3 FABSIX::R_GEEMon Nov 25 1996 19:085
	Field Service connected pump #3 into pump #2's outgoing slurry
	delivery line. This will allow DI water to flow to the secondary
	table. Calibrated all 3 pumps.

					A shift
21.125Polish table Motor and Gearbox replacement updateFABSIX::R_GEEMon Dec 02 1996 19:037
	
	All is on schedule, Field Service remove all obstructions.
	The old motor and gearbox are ready to be removed from the system. 
	The riggers will be in tomorrow morning to move the old motor 
	and gearbox out and put the new one in. Everything should be left 
	as is until morning.	
					A shift
21.126Polish table motor and gearbox replacement updateFABSIX::R_GEETue Dec 03 1996 19:338
	
	Installed new polish table motor and gear box. Field Service 
	re-installing everything that had been removed due to the 
	installation. They hope to test the table rotation before 
	leaving tonight. Facilities still needs to hook the exhaust
	back up. The system should be left as is for the night.
 
					A shift
21.127Slurry pump #2 problemsFABSIX::R_GEEWed Dec 04 1996 19:148
	The polish table motor and gear box has been replaced and 
	verified good, slurry pump #2 is stopping intermittently.
	As of the end of "A" shift pump #2 is still not functioning
	properly, field service is still working the issue. The
	system will remain down until the pump is fixed, read T.D.'s
	passdown for more info.

					A shift
21.128pump motor fixSUBPAC::CMPWed Dec 04 1996 20:365
    Replaced the analog output board because slurry pump 2 would
    intermitently run. when board was removed we found some of the jumpers
    to be misplaced. installed the new board with the correct jumper
    placement and tested. We would like to try the old board tomorrow
    to see if the jumpers were wrong or the board was actually bad.
21.129Pads changed due to failing MQC'sFABSIX::B_JUBINVILLEThu Dec 05 1996 05:456
    
           OPS couldn't bring uniformity in after a carrier change,
        off-set program and ID/OD.  Both pads were replaced and 
        weekly PM completed.
    
                                         D shift
21.130Wafer left on pad at lift-offFABSIX::R_GEETue Dec 10 1996 19:026
	The system ran good all day, at 6:00pm	a product wafer was 
	left on the pad at lift-off from carrier #2, last set off 
	wafers from a 30 wafer run. Re-seated the wafer and unloaded 
 	in manual. Had to reboot to clear the wafer tracking screen.
 
					A shift
21.131small slurry leak pump 1STRATA::FINNSun Dec 15 1996 18:327
Respond to moisture sensor alarm. Found it to be in the slurry pump 
cabinet. Leak was very small, located inside pump head on slurry 1. Repaired
leak and cleaned sensor and box. Tested aok.

						a shift

21.132? noisey bearingSUBPAC::CMPFri Dec 20 1996 01:038
    Received a call for an ungodly noise coming from the second table.
    Checked the flange at the pulley and found a pile of rusty metal dust
    and when the table was started there was a noise coming from that
    general area. Coordinating parts and Phx. support for problem.
    
    
    
                                                   SF
21.1332ND TABLE BEARING TRASHEDSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringFri Dec 20 1996 06:3416
         While running MQC's....   A loud continuous dragging,
      grinding squealing noise was coming from under/inside 
      the tool.  While running the tool in MANUAL mode with 
      side covers off, the source of the noise was definitely
      coming from the the 2nd table pulley assemble where the
      belt goes around.  A large pill of the wonder remains of
      bearing guts, I mean dust was all around the base of the
      shaft assembly.  We did the usual phone-a-rep and spoke
      with Rich.  He came in and evaluated the situation and
      got right on the horn with Pheonix and to get the ball
      rolling on the the dis-assembly and getting the proper
      parts sent in ASAP.  The carriers were pulled off the 
      tool and put in water and the pads pulled off.

      
                                             D shift
21.134second table bearing pulled parts on there wayFABSIX::J_SWENSONFri Dec 20 1996 19:304
    Continuing on with repair.
    Parts will be in by 2300.
    
                     C-shift
21.1352nd TableFABSIX::B_WESTPHALSat Dec 21 1996 19:4326
    
CMP.B3 - Down for a loud grinding noise coming from the 2nd table.  Closer
	 inspection revealed that the lower bearings were damaged. The lower 
	 bearings and center shaft for the 2nd table were removed.  The
	 bearings housings were rusted, some of the roller bearings themselves
	 were missing and The races were scored.  All the parts were ordered by
	 Speedfam but not all of them made it here.  We are still waiting on
	 some bolts and the lower shaft assembly.  Those parts should be here 
	 between 11:00 and 12:00 tonight via courier.  The reps should also be
	 here about the same time but they requested that we have someone at 
	 the desk to pick up the parts just in case.  There is some question 
	 about the 2nd table.  It was noticed that there is a little rocking 
	 motion in it and it rumbles when spun by hand.  Speedfam speculated
	 that this will stop when the shaft is reinstalled.  If that is not
	 true, Tim insists that the table be pulled and that upper bearing
	 inspected.  Tim also wants to insure that the bolts securing the shaft
	 to the top of the platen be sealed with RTV and given time to cure. 




	




21.136bearing replacementSUBPAC::CMPSun Dec 22 1996 03:167
    Replaced the lower pulley/bearing assy. for the second table, ran tool
    with test wafers and noticed no unusual noises or operating conditions.
    The tool is now quieter than a sleeping church mouse.
    
    Upon checking with manufacturing about the rumbling noise coming from
    the table when it was turned by hand, unless there is sticking or
    binding the bearing should be fine.
21.137new pad and 2, 3, and 5 carriersLUDWIG::RIDLONEliminate the obliviousTue Dec 24 1996 07:325
    Changed pad and carriers 2, 3, and 5 due to slow removal and high
    uniformity.
    
    
    						B-shift
21.138mapper timeoutSUBPAC::LANDRYTue Dec 24 1996 13:476
    The mapper failed today (timeout), the "B" mapper more specifically.
    Powered off the mapper controller for a couple minutes.  Repowered 
    said controller.  Re-homed the Load Station.  Good to go!
    
                                               A shift
    
21.139EP Regulators, Down Force Calibrated due to Removal & Uniformity ASDG::POIRIERThu Dec 26 1996 19:1928
    
    Head to Head Removal Rates & Head to Head Uniformity out of Spec:
    -----------------------------------------------------------------
    The head to head removal rates were > 900 A. The EP Down Force
    regulators were checked. EP #5 was 1. lb low at 59 lbs. (4095)
    All others were within .10 lbs. of 60.
    
    The down force was recalibrated, the results are as follows:
    
         Before                  After
    
    1)   377.6                   n/a
    2)   371.0                   376.5
    3)   389.6                   377.8
    4)   385.4                   376.5
    5)   364.7                   377.2
    
      
    Mqc's to commence.
     
    
    
    Head to Head Uniformity:
    ----------------------
    Carriers 2 & 5 were replaced due to high uniformity.
     
    
    
21.140Pads and carrier #2 replacedSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringFri Dec 27 1996 05:517

        The tool failed for deviation miserably on #2 and the other
     (4) heads were edge fast.  After the pads and carrier #2 were 
     replaced, the tool came in.

                                             D shift
21.141Unload cup issueSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringFri Dec 27 1996 07:4411

           The unload flipper failed to take the wafer out of
       unload cup #3 earlier in the morning for no apparent
       reason.  It ran fine after that until it just did it
       again.  It was discovered that the unload cups were
       all one location off of their correct holes.  After the
       cups were returned to their respective holes, the tool was
       turned back over to OPS. 

                                             D shift
21.142Pad change due to low R.R.SUBPAC::BJUBINVILLECMP-6 Equipment EngineeringSat Dec 28 1996 07:056

          24hr MQC's showed removal rate down near 2400.
         Pad change at 375 wafers.... yea! 

                                             D shift
21.143Head to Head Removal rate probs / Pads suspectASDG::POIRIERSat Dec 28 1996 19:3931
    
    Head to Head Removal Rates Out of Spec:
    --------------------------------------
    D Shift replaced the Pad just prior to shift change this morning.
    Head #4 was removing 1000 A to 2500 A more than the other 4 heads
    on a product lot. The carrier on head #4 was removed and measured,
    the exposure was .008 -.009". The carrier was replaced with a new one
    which measured .007 - .008. An MQC was run with the origional
    offsets unchanged, the range changed to 600 A with carrier #4
    having the lowest removal of the 5 heads.
    
    At this point the offsets were recalculated and the range dropped to
    198 A.
    
    FYI: This tool had its down force and EP regulators calibrated
         on Thursday.  The initial offset calculation spread was 
         close to 30 right after the calibration and again today.
    
         B1's down force and EP's were cal'd on Friday and the offset
         calculation was between -22 and +20. 
    
         We usually see the initial offset cal under 10 right after a 
         cal. We've also been seeing the head to head removal and
         uniformity all over the place during the week. We suspect
         the pads are now causing problems with the Speedfams. 
         There are 200 + pads in the stockroom, try different lot #'s
         if you get into trouble. 
    
       
                                                                C Shift.
    
21.144No title requiredFABSIX::M_RIDLONEliminate the obliviousSat Dec 28 1996 22:189
          
          For the sake of uniformity (kind of an oxymoron since
          all four Speedfams are different) the unload cups were
          put back where they belong. Load 1 goes under the load
          flipper and unload 1 goes under the unload flipper.
          Unload flipper to unload cup alignment was checked before
          and after the change and no difference was seen.

                                        
21.145carrier 5FABSIX::M_RIDLONEliminate the obliviousTue Dec 31 1996 07:149
    Carrier 5 misloaded a wafer causing a main vac low alarm. This happened
    last night as well and a cup to carrier alignment check was done which
    showed the alignment to be good. A cup to carrier alignment tonight
    had the same results. Checked the solenoid and found nothing wrong
    their either. Raised the plunger up a thousands and did several manual
    carrier 5 loads and the wafer was seated every time. 
    
    						
    						B-shift
21.146oem vac filtFABSIX::M_RIDLONEliminate the obliviousSun Jan 05 1997 03:134
    INstalled OEM vacuum filter and another gauge.
    
    				
    						B-shift
21.147Diamond ring changed!SUBPAC::LANDRYTue Jan 07 1997 16:0510
    Today........
    
    The diamond conditioning ring was changed out per Process Engineerings
    request.  The wafer count (MECH_ACP WAFER COUNT key in the Service
    screen) was 47337 at the time of the change.
    
                                          A shift
    
    
                                
21.148conditioner ring change procedure usedFABSIX::R_GEETue Jan 07 1997 17:1018
	The procedure used to breakin the new conditioner ring was taken
	from not 17.351. The procedure is as follows:
	
	- Did not remove old pad.
	- Checked pad step height to ensure there is >60 mils.
	- Removed old ring and labeled it as "Used ring from cmp.b3", it's
	  on the shelf behind cmp.a1
	- Installed New conditioner ring, sn#222.
	- Ran conditioner manually at 80% downforce for 30 minutes on old pad.
	- Changed pad.
	- Recorded total number of wafers on the tool(Note this is the total
	  running counter that is not re-settable). This is being done to 
	  track the total # of wafers on the conditioner ring.  Recorded this
	  number in notes.
	- Performed std 10 minute condition followed by normal breakins and
	  MQCs.

					A shift
21.149Moisture Sensor Alarm / Looking For CauseASDG::POIRIERThu Jan 09 1997 19:227
    
    Moisture Sensor Alarm:
    ---------------------
    The tool alarmed at 7:20 P.M. for a moisture Alarm. We are taking
    a look at the situation now. We'll have more info at passdown.
    
                                                               C Shift.
21.150 pad changeFABSIX::J_SWENSONSat Jan 11 1997 18:0014
    head to head variation out
    -------------------------
    
   primary pad    980 
    
   secondary      1000
    
   carriers       500
    
   primary pad thk     .070
    
    
    replaced primary and secondary pads
                                       C-Shift
21.151More carrier rinse problemsFABSIX::O_WHITFORDQuticherwhinin&#039;Sat Jan 18 1997 02:207
    Replaced solenoid valve #80 (carrier #2 rinse) as valve was sticking
    most of the time and spaying water all over creation & multihead. While
    finishing the monthly, carrier rinse #1 bagan to act up. Made minor
    adjustments to the pressure delivered to carrier #1 & #2 rinse as this
    seemed to be what was delaying or preventing the valves from closing on
    time. The spray still fans nicely, it just doesn't reach the ceiling
    anymore.... :p      
21.152CB5 trippingFABSIX::M_RIDLONEliminate the obliviousTue Jan 21 1997 03:2710
    Tracking a potential problem....on Saturday night CB-5 tripped after
    the operator checked the pad step height. Proper procedures were
    followed by the operator in that the OPEN PROCESS DOORS button was
    pressed. No adverse conditions resulted Saturday night but tonight it
    happened againg during the weekly and this time the infranor for the
    pad conditioner errored (L3) and we had to power down the tool to bring
    it back up. 
    
    
    					B-shift
21.153sticky fingersSUBPAC::CMPThu Jan 23 1997 16:045
    Intermitently dropping a wafer at unload flippers.
    Found that some water infiltrated the unload flipper box, cleaned
    the water and dirt, but found that the air cylinder was binding,
    not allowing the fingers to close completely. Replaced the cylinder
    and tested with no errors.
21.154Wafer not present errorsFABSIX::O_WHITFORDLead, follow or step aside.Thu Jan 23 1997 23:5416
    Wafer not present errors on carriers 1,2,3,& 5
    
    Check vacuum system starting with the filter. Vacuum on pump side was
    low at 20" Hg, even worse on tank side at 18" Hg. With wafers on the
    carriers and the MHTA at polish table, vacuum slowly dropped to about 
    12" Hg on the tank side. Replaced the filter element and got vacuum back 
    to 22" Hg on the pump side, about 20" Hg on the tank side, but it still 
    failed with the MHTA at polish.
    
    Checked vacuum line from tank to MHTA, reseated the lower connection from 
    the line leading thru the flex track, now vacuum on pump side is 26", and
    24 on the tank side.
    
    Ran 15 break in wafers with out a hitch.
    
                        Dave & Leon 
21.155System Pausing During Load / Scrubber Time out ASDG::POIRIERSat Jan 25 1997 10:5139
    
    Load Flipper Timeout Error: (C Shift)
    --------------------------
    Fri. 1/24/97  
    C Shift neglected to include this in the passdown last night.     
                                           
    The Load Flipper lowered to the Prealigner, closed onto a wafer and
    rose up approx 1 inch. The Flipper then stopped and the auto cycle
    paused. The Operator pressed START and the Auto Cycle resumed without
    incident.
    
    
                                                            C Shift.
    
  -----------------------------------------------------------------------------
    
    
    Load Flipper Timeout Error: (D Shift)
    -------------------------- 
    Sat. 1/25/97
    D Shift neglected to include this in the passdown this morning.
    
    The Load Flipper lowered to the Prealigner, closed on the wafer and   
    rose slightly. The Flipper stopped and the Auto Cycle paused.
    The operator resumed the cycle by pressing the START button.
    
    -------------------------------------------------------------------------
    
    Scrubber Timeout Error: (C Shift)
    ----------------------
    Sat. 1/25/97 10:30 a.m.
    
    An Assist was logged by the Manufactring Tech for thr Scrubber Timing
    out. The Scrubber moved forward, scrubbed the wafer and failed to 
    return to the home position.
    
    
                                                                  C Shift.
    
21.156mapper errorFABSIX::M_RIDLONEliminate the obliviousMon Jan 27 1997 20:535
    Tool alarmed for MAPPER 1 and MAPPER 2 errors during the start of
    process run. Homed the load module which corrected the problem and tool
    mapped as expected.
    
    						B-shift
21.157Software Utility Progran Run/No Corrupted FilesASDG::POIRIERThu Jan 30 1997 19:177
    
    Software Utility Program Run:
    ----------------------------
    Norton Utilities was run today by Speedfam. No corrupted files were
    found. 
    
                                                            C Shift/Dave.
21.158Pad change - scheduledSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringFri Jan 31 1997 07:509
    
         The step height was checked for the MQC's and
      it was around .51.  The wafer count was actually
      over 2000 too.  Both pads were replaced and the 
      primary table got (2) conditioning cycles.  It is
      WETMODE and needs full break-ins and MQC's.
    
                                       D shift
    
21.159MHTA/infranor and Mapper issuesSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringThu Feb 06 1997 06:1616
     The mapped failed to do its thing.  A hard boot didn't cure
   it.  At the the time of the failure, the other two Speedfams
   were not available so we manually mapped the boat to get them
   running.  When the OPS tech went to hit START on the product
   run.  The MHTA went to LOAD position but then timed out for
   NOT ARRIVING AT LOAD."  We went into MANUAL MODE and sent the
   MHTA to polish and then back to load again.  It timed out again...
   After this point, the MHTA wouldn't move at all.  We noticed
   there was no LED display on the infranor.  We tried hard-booting
   the tool to try and restore power to it but it was cooked.  
   We replaced the infranor and it worked fine.  The mapper was working
   fine as well...


                                             D shift
21.160wafer loss adjustFABSIX::M_RIDLONEliminate the obliviousSun Feb 09 1997 03:445
    Wafer loss detect in segement 1 of tupa recipe. Found 2 of the sensors
    needing to be tweaked. Resume processing with out error.
    
    
    						B-shift
21.161blah blahFABSIX::M_RIDLONEliminate the obliviousWed Feb 12 1997 06:006
    Pads and carriers changed due to engineering lot contamination. MQC's
    failed on head 4. It was changed again and given back.
    
    
    					  b-shift
    	
21.162B3 ?SUBPAC::CMPWed Feb 12 1997 19:494
        B3 lost a wafer after its run when the arms would lift off from the
    main p. table. No one was able to tell me for sure which one droped the
    wafer. I ran 50 wafers with the same recipe called "LIGHT". I had no
    further problems. System was put back up.
21.163N2 GUNFABSIX::B_WESTPHALSat Feb 22 1997 19:576
CMP.B3 - Installed an N2 gun on the system.


    


21.164FILTER HOUSINGS MOVED.FABSIX::B_WESTPHALSat Mar 01 1997 21:276
    
CMP.C1 - Moved the slurry housings from the front of the tool and put them in 
	 the slurry distribution box as engineering requested.  There are no
	 filters in the housings so they will need to be installed on Monday.


21.165door fix/upgradeSUBPAC::CMPMon Mar 03 1997 19:345
    Problems with doors. Found doors to be falling down with hardware
    loose.
    Replace door assy's with new style upgrade, wire new switches, and
    locks. Return tool to operation for verify and MQC's
     
21.166inverter errorFABSIX::R_POIRIERWed Mar 12 1997 07:2513


		While in wet idle, an Invertor Communication Error
		occured. Problem was one of the process doors wasn't
		making contact. Fixed the doors, and tried to run
		a lot and had Carrier #2 rotation error. Soft-
		booted tool and checked carrier rotation. OK
		Re-started lot and it is running fine.


					B Shift

21.167Vacuum errorSUBPAC::BJUBINVILLECMP Equipment EngineeringThu Mar 13 1997 05:4914

         The tool was alarming for NO WAFER PRESENT in #1,
      #3 and #4.  The OPS tech physically checked to see if
      they were there and in the pocket correctly, which they
      were.  He hit start but the wafer in one of the carriers
      slid out at touchdown and he EMO'ed the tool before 
      anything happened.  We checked some of the HEAD to CUP 
      alignments and they were ok.  We re-ran wafers to check
      the vacuum and everything was fine.  The tool came back
      up without any problems.  


                                             D shift
21.168Wafer Breakage / Root Cause UnknownASDG::POIRIERThu Mar 13 1997 17:2044
    
    Wafer Breakage:
    --------------
    Three wafers were broken and two scratched during a run this morning.
    
    The following adjustments were checked:
    
    EP Regulators: @ 4095
    
     Head #                  pre                             post
    ---------------    -------------------         ---------------------
         1                   59.6                          60.0
         2                   59.9                          60.0
         3                   59.8                          60.0
         4                   59.8                          60.0
         5                   59.6                          60.0
    
  -----------------------------------------------------------------------------
    
    Down Force:    Point to point method.
    
     Head #                  Pre                           Post
    -----------         --------------                ---------------
       1                   353.6                          n/a
       2                   348.2                          354.2
       3                   350.0                          354.0
       4                   353.2                          353.8
       5                   353.6                          n/a    
    
     Wafer loss detection sensor #'s 3 & 5 were readjusted. 
     
    
      The pad had 1000 wafers run and the carriers had 212.
    
      We're not sure what happened, root cause unknown.
    
    
      Later in the day, when the MQC"S were being run, carrier #3
      rotation started oscillating. DCX board #3 had error lights
      4 & 5 lit. After the DCX board was reset, all carriers 
      failed to rotate. A soft boot was then performed and the 
      problem disapeared. 
    
                                                       C Shift.
21.169Stall - Crash - Software CorruptionSUBPAC::BJUBINVILLECMP Equipment EngineeringSat Mar 15 1997 06:4020

          The tool had excessive vibrations during a touch-up
       run and the drive belt was starting to jump and snap.
       The tool was EMO'ed before anything crashed.  We checked
       the slurry flow and it was fine.  The TUP_B recipe was
       checked for any problems and it looked fine.   We started 
       up the tool with the same recipe to see if we could see a 
       problem.  While looking for a problem during the polish 
       step #5, a break-in wafer slid out of head #5 and crashed.  
       Dave from Speedfam was buzzed and he came in.  He was looking
       into when we found out that the crash Thurs. morning
       had the same symptoms.  We suspected corrupt software
       within TUP-B because we ran BIN2, BIN3, B-IN and TUP_A
       with no problems.  We ran TUP_B and it stalled.  We
       deleted TUP_B and copied over IO_MQC.  After that,
       TUP_B was ran and the tool was up and running. 


                                             D shift
21.170Wafer Loss Detection Board Replaced.ASDG::POIRIERSat Mar 15 1997 10:2016
    
    Wafer Loss Detection Board Failure:
    ----------------------------------
    In response to Thurday mornings loss of 4 product wafers, a test of
    the Wafer Loss Detection system was performed.. While running the NOTOUCH
    recipe, a wafer was placed in the beam path of wafer detect sensor #3. 
    The alarm sounded and an error message appeared on the crt. However,
    the contactor did not pull out and the polish table continued to 
    rotate.
    
    The wafer loss detection board was replaced and the above mentioned
    test was repeated. This time the contactor pulled out immeadiately
    and the table stopped rotating.  
    
                                                            C Shift.
    
21.171FABSIX::M_RIDLONEliminate the obliviousMon Mar 17 1997 05:5410
    Carriers 2 and 4 uniformity were running high out of spec. Changed them
    and zeroed the df offsets. MQC's were better but removal rates for
    carrier 4 are still running lower than the rest (same problem last
    week when the Monthly was started). Its not out of spec so production
    is still running. The air lines for carrier 4 df were checked in the
    multihead and no loose connections were found, plus no other
    abnormalities were seen either. 
    
    
    						b-shift
21.172Rotation #2 blown infranorSUBPAC::BJUBINVILLECMP Equipment EngineeringWed Mar 19 1997 06:2412

      Tool was alarming after loading (5) wafers for circuit
   breaker #5 tripped.  Came in a found breaker #5 tripped 
   so we reset it.  We then checked the infranors and found
   Rotation #2 dead.  #5 doesn't control Rotation #2 though.
   hmmmm....  Tried resetting tool to see if rotation #2 would
   come back but it wouldn't.  We ended up replacing the 
   infranor and bring the tool back up.


                                             D shift
21.173BROKEN SENSOR FABSIX::B_WESTPHALWed Mar 19 1997 18:339
    
CMP.B3 - Down for a" no slurry available" alarm.  Asland did a filter change 
	 which caused a bubble in the line.  When I moved the sensor on the
	 line to a place where slurry was, to clear the alarm, I accidentally 
	 broke the sensor.  Engineering is going to order a new part tomorrow.
         I temporarily jumped out the sensor until the new parts arrive.  The 
	 tool is presently up with a warning.


21.174replaced slurry sensorFABSIX::B_FINNWed Mar 26 1997 19:295
	Replaced broken slurry sensor. Cal/tested aok


						A shift
21.175wafer slide out segment 3 of tup_bFABSIX::J_SWENSONFri Mar 28 1997 19:2220
    tool crashed
   --------------
    
    
    - A wafer slid out of carrier four and into carrier 3 before the wafer
      detect system stopped.
    
    - Both wafers scrapped were product. (see note 9.55 for crash report)
    
    - The error happened on recipe tup_b and it was during segment 3.
    
    - The tool has been cleaned, carriers removed and pads removed.
    
    - most probable cause of the slip may be an intermittent loading
      issue. The tool ran fine all day.
    
    - will pass on to D-Shift
    
    
                                          C-Shift
21.176#4 arm problems source of crashSUBPAC::BJUBINVILLECMP Equipment EngineeringSat Mar 29 1997 04:1315

          Cleaned up tool after crash and started looking into 
     crash cause.  All of the alignments were checked.  Heads #2, 
     #4 and #5 were out.  Head #4, which was the carrier that the
     was the source of the crash, was out almost 3/8".  After all
     of the positions were corrected, we noticed that head #4 kept
     taking a long time to home.  When at home, the INPUT for #4
     AT HOME was almost blinking.  We ended up having to adjust the
     position of the home switch.  After that, we had to go through 
     with checking the LOAD position again.  The tool was turned back
     over to OPS for break-in for the new pads and carriers.  


                                             D shift
21.177Slurry Loss Detector brokeFABSIX::R_POIRIERTue Apr 01 1997 06:329
    While doing the Weekly PM on B3, I broke the Slurry Loss detector.
    I couldn't find a part number for it in SGMR, so I jumpered it out.
    I also put a warning on the tool for this problem.
    
    
                				B Shift
    
    
    		
21.178CMP.B3 crashFABSIX::J_SADINFreedom isn&#039;t free.Tue Apr 01 1997 20:3516
    
    
    	System errored for wafers not on carriers in segment 1. Checked
    carriers and found wafer in carrier 1 to not be seated properly. redid
    alignment on carrier one. All was ok until carrier three slid a wafer
    in subsequent runs. We lost carrier 2 and carrier 3, but managed to
    save 4,5,1. Alignment fixtures are mounted on system, but we haven't
    had a chance to align anything yet (multiple software glitches have
    required reboots galore). A new primary pad has been installed and the
    conditioner has been cleaned. Please reset the primary pad count when
    you have a second.
    
    	Carriers 1,4,5 are sitting on CMP.B4 and two new carriers are on a
    cart in the room next to CMP.B3. 
    
    	a-shift 
21.179Come on SUBPAC - This is the 2nd time you screwed up!!SUBPAC::BJUBINVILLECMP Equipment EngineeringFri Apr 04 1997 23:1724

      ok....  what going on....


            From Tuesday night....


              <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]CMP.NOTE;1 >>>
                                    -< CMP >-
================================================================================
Note 21.179                  CMP.B3 EE/PE PASSDOWNS                   179 of 179
SUBPAC::BJUBINVILLE "CMP Equipment Engineering"       4 lines   2-APR-1997 05:47
--------------------------------------------------------------------------------
    
    
           Took over from A shift with the alignments.  Nothing 
        was found out of the ordinary.  All head-to-cup and
        carrier to cups-up alignments were ok.  Installed two
        new carriers and two new pads.  Tool was turned over to 
        OPS for MQC's which passed.           

                                                 D shift
    
21.180Slurry sensor replacedSUBPAC::BJUBINVILLECMP Equipment EngineeringFri Apr 04 1997 23:187
    
    
        Replaced the slurry sensor and housing.  Alarm tested
      ok..   It it sensitive enough to where when you check the
      alarm, all you have to do is close the valve.
    
                                                    D shift
21.181FABSIX::B_FINNTue Apr 08 1997 20:4121
 Wafer crashed in receive cassette, Jim felt it was a misplaced cassette, 
clean and ran test wafers aok. 

System has low been running low removal rate for weeks. Pad is due for a change
due to end of life. The conditioner disk which is on the system now is from 
the aruga per ee last week. He wants it changed with one from stock. (sitting
 on stainless table in bag in CMP chase). 

First Bruce wants pad conditioner checked for correct downforce. See CMP
note 31.1 for procedure also in the maint spec.  After making any necessary
adjustments please Install the new conditioner disk. Break this in on the old
pad to remove any loose diamond per note 31.15. "conditioner disk breakin".

After note conditioner disk breakin has been completed the new pad can be 
installed.

Please leave the old cond. dist labeled for return to Tim Dyer.


							A shift
21.182Pad conditioner downforce calibrationFABSIX::A_QUILLENNobody will get away with anythingWed Apr 09 1997 07:508
    Performed pad conditioner downforce calibration (adjusted to 77lbs).
    Installed new conditioner disk. Old disk is labeled and is on the table
    next to the sink. Polish and index pads were changed. Initial MQC's
    failed. ID OD was tweeked and the next run is expected to pass. Results
    will be available by passdown.
    
                                          B Shift
    
21.183cmp.b3 conditioner experimentsFABSIX::J_SADINFreedom isn&#039;t free.Wed Apr 09 1997 20:4718
    
    
    	Checked pad conditioner downforce again...still at 77lbs (nice job
    b-shift). Checked conditioner centered on pad (good). Did slurry flow cal
    and found we were only flowing 400ml/min when calling for 500ml/min.
    Adjusted slurry flow until it matched what was called for. Mqcs were
    run again with results still being substandard (no improvement).
    Removed conditioner from CMP.B2 and placed on CMP.B3. Ran 300second
    conditioning recipe and then 5 breakins and 5 mqc wafers.  Results are
    pending (wafers still reading on UV-1050). Will enter results into the
    next note. 
    
    	CMP.B2's conditioner was returned to CMP.B2 and CMP.B3's original
    conditioner was put back on after mqcs were completed. 
    
    a-shift 
    
     
21.184cmp.b3 mqc resultsFABSIX::J_SADINFreedom isn&#039;t free.Wed Apr 09 1997 20:5713
    
    	Results for CMP.B3 with conditioning ring from CMP.B2
    
    
    	       mean	%dev	rate/min
    head 1 =   5586	8.5%	2793
    head 2 =   5631	8.1%	2815
    head 3 =   5610	8.6%	2805
    head 4 =   5381	6.4%	2690
    head 5 =   5850	8.1%	2925                         
    
    
    
21.185Tool has over 3000A R.R. with new pad....SUBPAC::BJUBINVILLECMP Equipment EngineeringThu Apr 10 1997 07:1915
       Tool was in spec yet the OPS tech logged it
   down  at shift change.  the tool was cleaned up 
   and while we had it, we did the WEEKLY PM.  We 
   put the same pad from the same lot that is on B2
   and the tool passed with over 3000A removal rate.
   B1 MQC's were run and that passed with the same numbers
   it has been for the last year...  Process Engineering 
   lifted the WARNING on the tool.

                                 Rock'n'roll baby!!!! 



                                             D shift
21.186Liquid detect unit/ InfranorSUBPAC::BJUBINVILLECMP Equipment EngineeringThu Apr 10 1997 08:4815
    
        On a side note...

        While doing the PM, the peristaltic pump box
    had slurry/water all inside with no LED on the
    leak sensor...  why???  Someone shut the detection
    unit off.

        The tool had been powered down before we were
    going through the containment pans and slurry lines
    being replaced when we powered up the tool and we
    apparently lost head #5 rotation Infranor.  It was 
    replaced and was good to go..

                                   BJ
21.187Slurry Detect Sensor, Fluid's cabinet leak, Slurry ysprayano carrier4FABSIX::J_SWENSONFri Apr 11 1997 15:5234
    Slurry detect sensor
   ______________________
    
            After filter change a bubble set off the sensor and would not
            reset.
    
            - checked 24 vdc npf
            - checked continuity between connector npf
            - Found switch setting to be off
            - corrected switch setting
    
            - verified sensor by removing from housing setting the sensor
              off
    
    
    
        Fluids Cabinet
        --------------
    
                 Found raw water leaking at several fittings, replumbed and
                 tightened, no leaks found.
    
                 Found cwr line to be leaking tightened and no leaks found.
    
    
     *** Noticed Slurry was spraying off of the madonna cup and hitting
         carrier #4. Bruce would like someone to look into the issue at
         hand. Tool had to be returned for production since B1 was down
         for the weekly.
    
    
    
                                           C-Shift
                 
21.188SLurry Detect ErrorASDG::POIRIERFri Apr 11 1997 20:2115
    
    Slurry Sensor:
    -------------
    The system alarmed for "NO Slurry" during the scheduled filter change.
    The alarm would not silence. The slurry detect module switch position
    was changed to the MIDDLE position and the problem ceased.
    
    
    LEAKS:
    -----
    Small DI and raw water leaks were repaired in the fluids cabinet.
    
    
    
                                                            C Shift.
21.189adjust temp.SUBPAC::CMPTue Apr 15 1997 20:084
    The Tool has had some problems of late with low removal rates, one
    thing that is being tried is to bring up the temp a couple of degrees.
    Worked with Tech to adjust the cdata.cmp for the chiller coeffs to
    bring temp. up to 26. Will monitor at new temp for any result.
21.190tempFABSIX::B_FINNTue Apr 15 1997 20:479

	Actual temp readings at polish table show table 4'c colder than set
 point. At EE request we adjusted input and output coff to raise temp Setpoint
 to 27'C. 

				A shift / Rich


21.191New pad also put on previous to temp adjustLUDWIG::BTOWERWed Apr 16 1997 08:223
    Before the temp was adjusted a new pad from the stockroom(newest
    delivery) was put on, broken in and MQC's run. Removal Rates are
    still averaging 2800-2900.
21.192Removal problemsLUDWIG::STANDINGWed Apr 16 1997 20:563
    CMP.B3  Head removal rate problems, head 2 was normal, other four heads
    were approx 2400 ang. Down force caled. head 2 replaced, runnin mqcs
    
21.193Low removal checks.LUDWIG::BTOWERFri Apr 18 1997 08:417
    Took slurry sample from system yesterday. % solids = 12.74. Specific
    gravity = 1.074. Also checked down force recipes on B1,2 and 3. Each
    system uses recipe 1(2 step recipe). All recipes matched except
    B1 had 77% down force in step 2 as opposed to 67% for the other two.
    Matt will be bumping up B3 today to see if it buys us anything with
    the low removal. Still awaiting word from speedfam on cut of
    conditioner. Reps also called them today for that info.
21.194TRANSFER PADDLE REPLACEDFABSIX::B_WESTPHALFri Apr 18 1997 21:476
 
CMP.B3- Down for problems with the unload transfer paddle.  The paddle seemed
	to be sticking when traveling toward the wafer tracks.  The problem 
	was the tol-a-matic (the assembly the paddle rides on).  The same was 
	replaced and adjusted.  Tool was returned to production. 

21.195Polish table belt and other problemsFABSIX::R_POIRIERMon Apr 21 1997 07:4433


		While running a process (Tup_B) the polish table
		began vibrating alot. We think the vibrations
		caused the table belt adjustment to loosen which
		started causing the banging noise that happens
		when the belt slips. We checked the belt and it
		was very loose as were the adjustment bolts. We
		re-tightened the belt and ran more wafers. There
		was still some vibrations, but they weren't as
		loud. The second set of wafers vibrated even less.
		Ran two more sets of wafers using TUP_H and there
		were no vibrations at all. Recommended to the OPs
		people to have Process check out TUP_B. After we
    		tightened belt, during boot up, unload transfer
    		paddle would not home. We had to re-teach it the
    		home and both track positions.


		Also, earlier in the shift the operators found two
		broken wafers in the left unload boat. Wafers in
		slots 1 and 13 or 14 were broke (clean break). Unable
		to determine the cause of this. Cleaned elevator and
		continued running tool. Problem did not re-occur.
		Later also in the left 	unload, wafers in slots 7 and 
		8 ended up cross-slotted. I think this was due to the 
		boat being bumped while trying to push a wafer into 
		the boat during the run.


						B Shift

21.196tup_a recipe flaws caused vibrationSUBPAC::LANDRYMon Apr 21 1997 18:5315
    A followup to the vibration problem exhibited on B shift last night...
    
    The recipe (TUP_B) was found to have two flaws.  The first concerned
    the ramp time for the table speed.  It was incorrectly set.  The table
    was trying to make set point immediately (30 or so RPM), hence causing
    the belt to jerk excessively.  The second program problem was centered
    around slurry flow prior to carrier touch down.  Not sure what step it
    was, but it was set for 1 secound, when it should have been 20.  Both
    steps were corrected, all is well with this problem.  
    
    The Opts tech also checked the TUP_A recipe and found a flaw with that
    one as well.  It had 230 secound pre-touchdown of slurry flow.  A tad
    bit much, yes/no?  Edited that one as well.  
    
    						A shift
21.197New pads and four carriersFABSIX::A_QUILLENNobody will get away with anythingTue Apr 22 1997 06:395
    System went down for a polish and index pad change. Carriers #1, 3, 4
    and 5 were also replaced. Following successful MQC's, the system was
    put up.
    
                                          B shift
21.198mapper errorFABSIX::J_SWENSONSat Apr 26 1997 19:5317
    mapper error
    ------------
    
    - Tool would not map
    
    - Found ionizer not working
    
    - powered down controller
    
    - replaced ionizer
    
    - powered up controller
    
    - no more mapper errors
    
    
                               C-Shift
21.199Mapper time-outSUBPAC::BJUBINVILLECMP Equipment EngineeringWed Apr 30 1997 07:372
    
        FYI -   Two mapping time-out errors this morning...
21.200Mapper errorsFABSIX::R_POIRIERMon May 05 1997 06:1414
    
    		B3 had more mapper errors last nite. First,
    		the B side wouldn't map at all., tool thought
    		the wafers were cross-slotted. A side mapped
    		but the robot had problems picking up the wafers.
    	        We re-booted the tool and turned up the gain on
    		the ionizer. Ran wafers on both A and B sides
    		and had no more errors.
    
    
    
    					B Shift
    
    		
21.201Pad dueSUBPAC::BJUBINVILLECMP Equipment EngineeringSat May 10 1997 09:126
    
         Tool failed 24 hrs MQC's for 2396 removal rate average.
    
         Will pass to C shift for pad replacement.
    
                                             BJ
21.202CMP.B3 crash recoveryFABSIX::J_SADINFreedom isn&#039;t free.Sun May 11 1997 11:2512
    
    
    	Cleaned up 2nd crash on CMP.B3 (tool had crashed earlier on
    B-shift). Replaced both pads and all carriers. Checked head to cup
    alignment and height, as well as vacuum level and recipes. All appeared
    ok except for a time difference in step one of TUP-C (it was set to
    2secs instead of 25secs). Changed the time back to 25 secs. The machine
    had crashed during step 5 of TUP-C. 
    
    	a-shift
    
    
21.203safety upgradeSUBPAC::CMPTue May 13 1997 20:234
    Today we installed a MHTA safety stop. What this does is keep the
    multi-head lock from disengaging. To use it, requires the wedge shaped
    piece to be placed under the MHTA lock lever. This will prevent the
    multi-head from moving.
21.204CMP.B3 particle issueFABSIX::J_SADINFreedom isn&#039;t free.Tue May 13 1997 20:3121
    
    
         Operations alerted us to a high particle issue on this tool today
    (after 24hr mqcs). Both large and total particle counts were out of
    spec (filter change was done this morning before mqcs were run). We
    changed the primary and secondary pads and cleaned the tool thoroughly.
    After breakins and mqcs, particles were still out of spec for large and
    total. As luck would have it, there was a barrel of SC112 in the chase.
    We hooked the barrel up to CMP.B3 and ran another set of mqcs with the
    barrel slurry. Large and total particles were now well within spec
    (large < 50, total < 150). 

    	All "B" machines have been displaying higher-than-normal particles over
    the past two days. Ashland has been contacted and advised that we MAY
    be requesting a filter change this evening (this needs to be
    coordinated with ops since some tools are running now). 
    
    
    	in peace and light,
    
    	TEAM-A
21.205Particles.... GONE!!SUBPAC::BJUBINVILLECMP Equipment EngineeringWed May 14 1997 07:147
    
            Tool failed once more for particles just into our
         shift.  Contacted Ashland about another Slurry filter
         change.  After they swapped/replaced filters, B3 passed
         particles.
    
                                               D shift
21.206Wafer crash....SUBPAC::BJUBINVILLECMP Equipment EngineeringThu May 15 1997 09:2511
    
            Head #4 never blew off its wafer that was just polished
        and it was able to pick the next wafer by shear surface tension.
        The heads touched down with 2 wafers on #4.  The wafers spun
        on each other then the bottom wafer slid out and hit #3 before 
        wafer detect tripped.  Carrier issue??  Blow off solenoid?  Tool
        was 95% cleaned up.  C shift took over.  
    
           total loss - 1 pad, 2 carriers and 3 product wafers
      
                                                  D shift
21.207Slurry alarmSUBPAC::BJUBINVILLECMP Equipment EngineeringFri May 16 1997 04:066
    
         Tool didn't alarm for slurry loss.  No product was
      lost.  After slurry system cmae back on-line, the slurry
      sensor was adjusted and verified.
    
                                              BJ
21.2085/15 carrier 4 issueFABSIX::J_SWENSONFri May 16 1997 17:5817
    Carrier four wafer breakage for 5/15/97.
    ----------------------------------------
    
       This should have been posted yesterday.
    
       - Inspected and verified the the electrical pneumatic functions of
         carrier 4 and found everything to be functioning properly.
    
       - Tried to force di through the carrier and none would come out.
       - Tried to force cda through the carrier and none would come out.
    
       - Dissassembled carrier and tried the above and same results.
         The plate of the carrier seemed to be the culprit (Being clogged).
    
    
    
                          C-Shift
21.209#5 POOR UNIFORMITYFABSIX::B_WESTPHALBradley BP#661 EXT# 6610Sat May 17 1997 20:523
CMP.B3 - Had poor uniformity on carrier #5.  The same was replaced.

21.210Down force calibratedFABSIX::R_GEETue May 20 1997 20:1515
	Operations asked us to check carrier #2 downforce due
	to high removal. Using load cell # 511654 we found the
	downforce at 408 lbs. Checked heads #3 & 4, these were
	both way high also. Due to the high downforce found we decided 
	to remove the pads (count 1450) and due all 5 heads, all
	5 heads were out. After the downforce was completed we checked 
	head #1 using load cell #477411, it read 311 lbs, should have 
	been 354 lbs. If the downforce was done last using this load
	cell this would account for the high reading. Re-calibrated all 
	5 heads using the strasbaugh load cell witch was reading a happy 
	medium. We will bring both speedfam load cells to the cal. lab.
	tomorrow AM. and ask for a rush.
    
					A shift
                           
21.211Spill...SUBPAC::BJUBINVILLECMP Equipment EngineeringThu May 22 1997 05:587
    
    
             We ph tested the 3-5 gallons of liquid under B3 and
        it was DI/glycol.  we put 6-8 pillows under there to try
        and soak it up.
    
                                             D shift
21.212Chiller - Slurry supplySUBPAC::BJUBINVILLECMP Equipment EngineeringFri May 23 1997 06:1620

          The tool went through all of the break-ins fine
       and the 1st run of MQC's showed the removal rate
       average around 3200 but the range was 360, spec is 300.
       OPS ran the off-set program and the rates dropped almost
       500A a minute, failing bad!!!  We contacted Tim and we
       started to pull the chiller.  All of the fittings on the
       chillers have unions but they are all hard plumber.  Worst
       case, I needed to hack-saw one of the lines.  Leon was 
       working on the weekly PM when I had the hack-saw in my 
       hand when he indicated that Slurry #1 supply valve was
       turned off.  The slurry sensor was not sensitive enough 
       either so that had to be tweaked, so, all of the tests 
       needed to be re-done.  The chiller was hooked back-up 
       and filled.  When it up to 27 degrees, OPS started 
       running wafers again.  Once we know where we stand 
       with the MQC's, we'll go from there. 

                                             D shift
21.213GOOD TO GO!!!!!!SUBPAC::BJUBINVILLECMP Equipment EngineeringFri May 23 1997 07:1713
    
    
            	Update:   as a result of Leon's keen eyes and
            alertivness, he noticed that the slurry #1 supply
            valve was closed.  As a result, the chiller was not
            replaced and the original chiller was back on line.
            The current chiller flow is 2 GPM, temp at 27 degrees
            and the MQC's wafers passed unif. and the removal
            rate average waws 3400 with a range of 155.
    
                       the anchovie pizza paid off for Leon!!!! 
    
                                               D shift
21.214Load problemFABSIX::J_SWENSONFri May 23 1997 20:3423
    INVERTER ERROR
    --------------
    
                          Tool errored for wafer not in pocket
                          Tool locked up
                          Rebooted tool
                          Tool still locked up
                          found m2 not being pulled in.
                          24v interlock problem
                          ohmed out circuit o.k.
                          still timeing out
                          went door to door checking switches
                          next thing you know tool powers up no problem
                          went to duplicate original error
                          created wafer not in pocket error.
    
                          Tool will need to have the load cup alignment's 
                          checked.                   
    
                            Will pass on to B-shift
              
      
                                              C-Shift
21.215No load problem but...FABSIX::L_SHKOLNIKOVSat May 24 1997 08:2614
      The tool went down for vacuum or misload on carr#2. We checked carrier 
    to load cup alignment  - fine. Released tool to production. Ran set of
    MQC wafers with average ~2000� - too bad. So verified conditioner
    downforce, slurry flow looks fine. There is only one related problem
    that was found - the table rinse flow control valves were open all the
    way out. We closed them to match B2 setup. Pad and carriers were
    replaced. Two out of five carriers had polished guide rings. So
    another set MQC has been run, the tool came up fine and is running 
    product now.  
    
    
    
                                                      D-shift.
  
21.216C shift passdown regarding particlesFABSIX::B_JUBINVILLEWed May 28 1997 22:1317
From:	ASDG::POIRIER      "FAB6 CMP EQUIPMENT TECHNICIAN. PAGER # 254 FUNCTION 1. TEL:6610,5804,4774,6609." 28-MAY-1997 19:29:29.78
To:	FABSIX::B_JUBINVILLE,FABSIX::L_SHKOLNIKOV
CC:	POIRIER
Subj:	passdown



   Brian & Leon,

   Here is the latest situation:

 
   CMP.B3 - DOWN/UNSCHEDULED: High Particles. Brad has verified that the house
   slurry is not the cause. Try cleaning tool.

 
                                                            C Shift. 
21.217Particles gone....SUBPAC::BJUBINVILLECMP Equipment EngineeringThu May 29 1997 03:3418

           Came in to the tool being down for particles.  Cleaned
       entire tool.  Found brownish tint to the DI water around
       polish table area.  Found metal reflective tape on all of the 
       carriers and polish table drain tubes completely oxidized.
       This tape was used 5/22 during the day for the RPM testing.
       Removed all corroded tape and cleaned tool.  Inspected scrub
       brushes and all of the top brushes had the same color as the
       polish area.  Running NO_TOUCH showed that the brushes were
       causing adders so they were replaced.  An allen wrench,
       found in the scrubber tank and was causing the scrubber to 
       bind up, was removed.  It saw it's better days....  OPS
       ran particles and all heads passed and were all under 200 
       particles.


                                             D shift
21.218Multiple ErrorsFABSIX::J_SWENSONFri May 30 1997 20:3626
    Multiple errors
    ---------------
    
    
                    Tool errored for the following multiple times
    
                        - wafer in water track with no wafer being present
    
                        - pad conditioner not @ home
    
                        - robot errors
    
                        - lo di flow
    
    
    No physical problems seen
    
    Reset CPU, checked all cables
    
    All seemed to be O.K.
    
    Ran out of time will pass on to D Shift
                           
    
                                             
                                      C-shift
21.219Mappers/Tilts - D.O.A.SUBPAC::BJUBINVILLECMP Equipment EngineeringSat May 31 1997 07:4424

       Came with tool down with software errors for the robot
    conditioner and mapper.  We rebooted the tool when we got
    in and all we got was an error for the mappers not initializing.
    Tilt A was doing some funky moving and Tilt B was dead.  We
    swapped cables and the A was was dead and B was all screwed
    up.  Per Rich, we replaced the controller and everything was 
    dead.  We then replaced both tilts and the cables to keep
    everything as a matched set.  The tilts and mappers initialized
    fine but the signal never got back to the computer that they
    initialized.  All of the cables where checked and nothing was
    found.  Rich feels we had a bad controller in stock.  He is
    getting in another complete Mapping system in today.  The
    tilts in the tool were shimmed just to get the tool running 
    in manual mode.  The mapper is powered off and is disabled
    in HARDWARE.CMP.  There is a WARNING on the tool for OPS to
    manually map the cassettes and not to use AUTO mapping.   
    Rich will be calling in the a.m. with some shipment updates.

                                           D shift 
  

                                             D shift
21.220Wafer crash in scrub stationSUBPAC::BJUBINVILLECMP Equipment EngineeringSat May 31 1997 09:0817
    
           Tool was running fine until ~07:30.  It appears that
       the top middle scrub brush fell off its shaft and landed
       on the bottom brush.  When the scrubber went in to scrub
       the 1st wafer of the bunch and BAM! CRASH! EXPLOSION!  The
       product wafer was in 1000 pieces.  The brush was floating 
       in track #1, the pedestal fixtures were everywhere.  The 
       next wafer was placed on one pedestal and fell to the 
       bottom of the scrub station.  One wafer wafer was scrapped,
       the one that fell and the one that was in the unload flipper
       need to be inspected by PSS.  They look ok...  unfortunately
       I have a weight problem and I can't get in to the scrub station.
       C-shift will lower Hawk by his ankles to clean out the debris. 
       The paddle might be bent and will need to be looked at as
       well.
     
                                                  D shift
21.221Mapper ProblemsASDG::POIRIERSat May 31 1997 21:1742
    
    Scrubber Brush:
    --------------
    The scrubber brush that fell off an D shift last night was replaced.
    The o-ring that belongs on the brush mount was missing.
    
    
    Contamination in Chiller lines:
    ------------------------------
    The chiller lines, platten water jacket & chiller tank were thoroughly 
    cleaned and flushed today. The chiller tank (screen/filter) element and
    regulator screen were clogged with debris. The flow meter sight glass
    was also disassembled and cleaned. It is now possible to accurately
    read the clilled water flow through the platen. With the chiller
    recirculation pressure @ 9 psi, the flow meter reads 2 1/2 gal/min.
    
    
    Mapper Problems:
    ---------------
    We followed up from D Shift on this problem ( See previous note). The
    new kensington mapper controller and tilt elevators were installed.
    When we placed a cassette on the tilt platform, it became apparent
    that the tilt platforms were for (6") cassettes! We removed the 6"
    tilt platforms and put the old 8" type & matching controller back
    in the tool. 
    
    The only snag we're facing at this time is the shims. (See previous
    note from D shift about shims under the tilt platform) The person
    who removed the shims has left for the day. We're not absolutly
    sure if they are back in as they were or why they are under the 
    platforms in the first place. B Shift, please verify that the robot 
    picking wafers out of cassette A & B properly before WUP'ing the
    tool. A robot adjustment may be necessary. The system will have to
    stay UP with a warning to MANUALLY MAP THE CASSETTES. 
                                                                  
    P.S. Rich Barrett from Speedfam was called about the mix-up and he 
    will send 8" tilt platforms and a matching controller ASAP.
    
                                                          C Shift.
    
    
    
21.222cmp.b3 chiller heater was offFABSIX::J_SADINFreedom isn&#039;t free.Mon Jun 02 1997 10:178
    
    
    	found chiller heater was turned off. This prevented the chiller
    from reaching its setpoint of 28�C (per Bruce Tower). Please leave the
    heaters on. Thanks.
    
    
    a-shift
21.223Moisture detection / Pad Cond InfranorSUBPAC::LANDRYMon Jun 02 1997 20:2325
Rough ride with this system today.......

The system initially alarmed for a "Moisture Detection" alarm.  Upon 
checking for the leak, it was discovered that there was NO leak.  All
sensors/papers were found to be dry and intact.  Hmmmmmm!  Acknowledged 
the alarm and alas, we were back in business.  About 20 minutes later,
the moisture alarm re-appeared.  Again, NO sensors were wet.  We allowed
the currently lot of wafers to be completed, then we took the system
down for a "hard boot".  Upon reboot, the system failed initialization
badly.  The Index table didn't move and the pad conditioner would go down.
We decided to reboot again.  The pad conditioners Infranor had no power.
Said Infranor was replaced and a reboot went fine.  The system then ran
for about an hour.  Low and behold, the moisture detection re-appeared.
We checked the control unit of the detection system and found the "leak
LED" began to flicker.  Wiring was checked.  A very loose wire was found
on TB 11.  This was tightened, but no luck.  

The control unit appears to be bad at this time.  It seems that it can't
compliment all the sensors.  We can unplug each one individually, and the
LED flickering stops each and every time.  We were going to change out 
the control unit, but none were in stock. SF representatives will try to
get one in for us ASAP.  At this time, the system is running with the
moisture detection shut off and under a warning state.

                                       A shift
21.224Mapper installed / no moisture detectSUBPAC::LANDRYWed Jun 04 1997 18:1321
    The mapper situation was finally addressed today.  An abundance of lots
    in the room early in the week was the cause for the delay in fixing
    this unit.
    
    A new mapper setup was installed.  This included the tilts, cables and
    the controller.  The installation went ok, with some calibration
    required post installation.  NO shims were used!  Thr mapper tilts were
    calibrated first.  Then the robot pickup positions were done on both
    tilts.  Finally, the robot to alignment fixture calibrations was
    performed.  We then cycled a multitude of wafers (NO_TOUCH) to verify
    the handling capabilities.  All went smooth.
    
    The original pad was salvaged as well as the carriers.  MQC's passed.
    The tool is cranking them out again.
    
    The tool is still running under one warning at this time.  The moisture
    detection unit did not arrive today.  The moisture detection circuit
    remains off line at this time.  The unit is powered off.
    
                             SF Representatives and the A shift
                                                            
21.225Brush fell off...SUBPAC::BJUBINVILLECMP Equipment EngineeringThu Jun 05 1997 07:118
    
    
          The top middle brush fell off and prevented a wafer
      to be scrubbed.  The brush was re-installed and it is a 
      real tight fit so I can't see how it came off.  Seems ok
      now.
    
                                         D shift